### v1

# CodePMP: Scalable Preference Model Pretraining for Large Language Model Reasoning

Abstract
Large language models (LLMs) have made significant progress in natural language understanding and generation, driven by scalable pretraining and advanced finetuning. However, enhancing reasoning abilities in LLMs, particularly via reinforcement learning from human feedback (RLHF), remains challenging due to the scarcity of high-quality preference data, which is labor-intensive to annotate and crucial for reward model (RM) finetuning. To alleviate this issue, we introduce CodePMP, a scalable preference model pretraining (PMP) pipeline that utilizes a large corpus of synthesized code-preference pairs from publicly available high-quality source code. CodePMP improves RM finetuning efficiency by pretraining preference models on large-scale synthesized code-preference pairs. We evaluate CodePMP on mathematical reasoning tasks (GSM8K, MATH) and logical reasoning tasks (ReClor, LogiQA2.0), consistently showing significant improvements in reasoning performance of LLMs and highlighting the importance of scalable preference model pretraining for efficient reward modeling.
1 Introduction
Figure 1: Compared to directly finetuning reward models, CodePMP significantly improves the sample efficiency and capability of reward models, which in turn boosts the generator’s(MetaMath-Mistral-7B) reasoning performance (Best-of-N accuracy) across both mathematical reasoning tasks (GSM8K and MATH) and logical reasoning tasks (ReClor and LogiQA2.0).
Large language models (LLMs) have made remarkable progress in natural language understanding and generation, benefiting from scalable pretraining and finetuning techniques like supervised finetuning (SFT) (Wang et al., 2022; 2023a) and Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022a; Lightman et al., 2023b; Bai et al., 2022b; Gulcehre et al., 2023; Schulman et al., 2017; Rafailov et al., 2024). However, enhancing LLMs’ reasoning abilities, particularly in complex logical and mathematical tasks, remains a significant challenge (Wang et al., 2023b; Zhang et al., 2024b). Although RLHF is effective, it relies heavily on high-quality preference data, which is costly and labor-intensive to annotate (Cobbe et al., 2021b; Zheng et al., 2024). This limitation impedes the scalability of reward model (RM) finetuning, which is essential for guiding LLMs toward optimal outputs.
To alleviate this issue, prior works like Anthropic’s Preference Model Pretraining (PMP) (Askell et al., 2021) have proposed improving reward modeling data efficiency by pretraining preference models on large-scale preference data from public sources like Reddit and Wikipedia, followed by an efficient finetuning on limited high-quality human-annotated data. However, this approach is less effective for reasoning tasks due to the scarcity of reasoning preference pairs available online. Compared to other tasks, manually annotating preference data for reasoning is inherently more challenging and difficult to scale (Zhang et al., 2024b; Zhou et al., 2023), highlighting the urgent need for a scalable PMP approach for reasoning tasks.
Figure 2: An illustration of the CodePMP process. First, raw code collected from GitHub is cleaned and summarized into code prompts (descriptions). Then, for each code prompt, a weak CodeLLM generates a rejected response, while a stronger CodeLLM produces a chosen response. Finally, these <chosen, rejected> pairs, accumulated in the millions, form the pretraining dataset for the preference model. This pretraining process improves not only sample efficiency but also the performance for downstream reasoning reward model finetuning.
In this paper, we propose CodePMP, a scalable preference model pretraining pipeline that enhances LLM reasoning abilities using synthesized preference pairs derived from high-quality, publicly available source code. Code, with its inherently logical and structured nature, provides rich data suitable for reasoning tasks. Recent works (Zhang et al., 2024b; Aryabumi et al., 2024) also show a strong correlation between code training and reasoning improvements in LLMs. By leveraging the huge amount and diverse coverage of source code available on platforms like GitHub, CodePMP offers a scalable solution for pretraining preference models, thereby improving RM finetuning efficiency and enhancing LLMs reasoning performance.
Specifically, CodePMP generates preference pairs by synthesizing chosen and rejected code responses for a given code-related prompt or description using CodeLLMs. A strong CodeLLM produces higher-quality (chosen) responses, while a weaker model generates suboptimal or even low-quality (rejected) responses. These <chosen, rejected> pairs, accumulated in the millions, form a large-scale synthesized preference dataset. This dataset is then used to pretrain the preference model with pairwise ranking objectives (Cobbe et al., 2021b; Charniak & Johnson, 2005), providing an good initialization for further fine-tuning the reward models.
We evaluate CodePMP on widely studied reasoning tasks, including mathematical reasoning tasks such as GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021), as well as logical reasoning tasks like ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023). Our experiments demonstrate that CodePMP significantly improves RM fine-tuning accuracy and Best-of-N performance in reasoning tasks, outperforming direct RM fine-tuning, as highlighted in Figure 1. Moreover, additional experimental results reveals that RMs initialized with CodePMP exhibit greater robustness across different tasks. These results indicate that code-derived preference data provides a scalable, cost-effective solution for enhancing LLM reasoning capabilities while reducing reliance on extensive preference annotation, achieving more effective reward modeling for reasoning tasks.
In summary, our main contributions are:
  1. 1.
    We introduce CodePMP, a scalable method that uses code-derived preference pairs to pretrain preference models, improving sample efficiency and robustness for downstream RM finetuning.
  2. 2.
    We validate that CodePMP significantly improves performance on reasoning tasks, demonstrating that a scalable PMP process positively impacts LLM reasoning abilities.
  3. 3.
    We provide a detailed analysis of key design elements in CodePMP, offering valuable insights for future research in related areas.
2 Preliminaries
Language Modeling
Language modeling (LM) is a fundamental task in natural language processing, aimed at modeling sequences of language. This is typically achieved through Causal Language Models (Causal LM), where the model is trained to maximize the likelihood of predicting the next word $w_{t}$ given the preceding words $w_{1},w_{2},\dots,w_{t-1}$. The training process minimizes the negative log-likelihood of the predicted word sequence:
$\mathcal{L}_{\text{LM}}=-\sum_{t=1}^{T}\log P(w_{t}|w_{1},w_{2},\dots,w_{t-1})$
This loss function $\mathcal{L}_{\text{LM}}$ encourages the model to capture the underlying patterns in the data. Transformer architectures (Vaswani, 2017) have become the standard for Causal LM due to their ability to handle long-range dependencies effectively.
Reward Modeling
Reward modeling (RM) is crucial in reinforcement learning from human feedback (RLHF), providing a scalar reward signal that guides the learning process based on output quality. The reward model $R_{\theta}$ predicts the quality of an output $y$ given a context $x$ as $s=R_{\theta}(x,y)$, where $s$ is the scalar reward score. In preference modeling, RM predicts the relative quality of outputs by comparing pairs. A common method is the Pairwise Ranking Loss, where the model assigns higher scores to preferred (chosen) outputs:
$\mathcal{L}_{\text{RM}}=-\log\left(\sigma(s_{\text{chosen}}-s_{\text{rejected}})\right)$
, where $s_{\text{chosen}}=R_{\theta}(x,y_{\text{chosen}})$ and $s_{\text{rejected}}=R_{\theta}(x,y_{\text{rejected}})$, and $\sigma(\cdot)$ is the sigmoid function.
Best-of-N Sampling
Best-of-N (BoN) sampling improves LLM reasoning (Cobbe et al., 2021b; Lightman et al., 2023b). In this approach, $N$ candidate solutions $\{y_{1},y_{2},\dots,y_{N}\}$ are generated by sampling from the LLM’s output distribution for a given problem. A reward model scores each candidate and selects the highest-scoring one as the final answer:
$\hat{y}=\arg\max_{y_{i}\in\{y_{1},y_{2},\dots,y_{N}\}}R_{\theta}(x,y_{i})$
, where $R_{\theta}(x,y_{i})$ represents the reward score for each candidate $y_{i}$. This technique is especially effective in tasks like mathematical problem-solving and logical inference, where selecting the most plausible solution from a diverse set of outputs improves overall accuracy (Wang et al., 2022).
3 Code Preference Model Pretraining
3.1 Model Design
Code Preference Model Pretraining (CodePMP) is designed to enhance the sample efficiency of reward models, particularly for reasoning tasks where high-quality preference data is scarce. Traditionally, reward models are finetuned on small, curated datasets, which limits their effectiveness in complex tasks like mathematical reasoning or logical deduction. CodePMP mitigates this limitation by introducing a pretraining phase between basic language model pretraining and finetuning on domain-specific reasoning datasets. This phase leverages a large, diverse dataset of code-preference pairs, enabling the model to learn generalizable patterns and ranking strategies.
CodePMP training involves two key components: Reward Modeling (RM) and Language Modeling (LM). In RM, the model is trained on code-preference pairs, learning to assign higher scores to the chosen code through a pairwise ranking loss. In LM, only the chosen code is used for autoregressive training to maintain the model’s general capabilities. The overall loss is a combination of the RM and LM losses, ensuring the model enhances its ranking ability without sacrificing general language modeling performance: $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{LM}}$.
3.2 Data Construction
To enable scalable preference model pretraining, we construct a dataset sourced from GitHub, which includes a diverse range of repositories and associated metadata. The dataset consists of two primary components: Repository Data comprises over 1.3 billion code files from GitHub repositories, while GitHub Metadata includes information such as commit histories, discussions, pull requests, and issues.
The CodePMP dataset is constructed through a systematic process. First, raw source code is processed by a description summarizer, typically an instruction-tuned CodeLLM, to generate prompts that describe the functionality of the code.
These prompts are then used by two CodeLLMs of different capabilities to generate code snippets:
  • Chosen response: Generated by a more advanced CodeLLM (e.g., 6.7B parameters).
  • Rejected response: Generated by a less capable CodeLLM (e.g., 1.3B parameters).
This process yields pairs of code responses—one chosen and one rejected—which are used for preference modeling. This scalable approach significantly enhances pretraining efficiency, improving performance on downstream tasks.
The steps of the CodePMP methodology are outlined systematically in Algorithm 1.
Algorithm 1 Code Preference Model Pretraining (CodePMP)
1:Source code repository $S$, Strong CodeLLM $M_{\text{strong}}$, Weak CodeLLM $M_{\text{weak}}$
2:Pretrained Model
3:Input: Source code $S$
4:Summarize description $D$ using $M_{\text{strong}}$ on $S$
5:for each $D_{i}\in D$ do
6:     Generate Chosen Response using $M_{\text{strong}}$
7:     Generate Rejected Response using $M_{\text{weak}}$
8:end for
9:Calculate LM Loss $\mathcal{L}_{\text{LM}}$ on Response
10:Calculate RM Loss $\mathcal{L}_{\text{rank}}$ using Chosen Response and Rejected Response
11:Train PMP Model using $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{LM}}$
4 Experimental
In this section, we first outline the experimental setup, followed by the experimental results, highlighting that CodePMP is a highly scalable method.
4.1 Experimental Settings
4.1.1 CodePMP Settings
Data Construction
We generate code preference pairs following Algorithm 1, using the deepseek-coder-6.7b-instruct model as the strong CodeLLM to generate chosen responses and the deepseek-coder-1.3b-instruct model as the weak CodeLLM to generate rejected responses. The constructed CodePMP dataset includes 28 million files and 19 billion tokens. The diverse datasets provide sufficiently broad prompt coverage for preference model pretraining, which is conducive to the generalization of preference models in reasoning tasks. In addition, the average lengths of the chosen and rejected responses are similar, ensuring that response length does not bias the CodePMP learning process. Details are provided in Appendix C and Table 4.
CodePMP Training
By default, we initialize the preference models with the publicly available Qwen models (Yang et al., 2024a), using different model sizes, specifically Qwen2-1.5B and Qwen2-7B. Detailed hyperparameters for CodePMP training are provided in Appendix B.
4.1.2 Reasoning Finetuning Settings
We validate CodePMP on reward models across two reasoning task types: mathematical and logical reasoning. The reward model is finetuned on corresponding preference datasets for each task. For mathematical reasoning, we use the MathShepherd-pair dataset, derived from MathShepherd (Wang et al., 2023b), and evaluate the model on a holdout test set to assess RM accuracy.
Similarly, for logical reasoning, we use the ReClor-pair and LogiQA2.0-pair datasets, derived from ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023), respectively. We train reward models on these datasets, with holdout test sets used to evaluate model accuracy. Dataset construction and finetuning hyperparameters are provided in Appendix D and  B.
4.1.3 Evaluation Settings
Following (Zhang et al., 2024a), we adopt two evaluation metrics:
RM Accuracy
This metric measures the accuracy of the reward model in distinguishing chosen from rejected solutions on the holdout test sets. It provides insight into the model’s ability to classify individual sequences.
Best-of-N (BoN) Accuracy
This metric evaluates the proportion of correct solutions selected by the finetuned RM from $N$ candidate responses. It assesses the model’s group-wise ranking performance, focusing on its ability to select the correct answer from a set of candidates. We use MetaMath-Mistral-7B (Yu et al., 2023) as the generator for this evaluation.
For mathematical reasoning, we use the GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021) test sets. For logical reasoning, we evaluate on the ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023) test sets. Further details can be found in Appendix D.
Note that logical reasoning questions typically involve a paragraph followed by statements to be judged true or false, making Best-of-N evaluation challenging. Therefore, we use multiple-choice accuracy, where the reward model ranks four manually annotated options and selects the best one. This metric is equivalent to Best-of-4, and thus, for logical reasoning tasks, multiple-choice accuracy and Best-of-N are used interchangeably.
4.2 Experimental Results
4.2.1 RM Accuracy Results
We first compare RM accuracy on the holdout test set with and without CodePMP initialization. As shown in Table 1, RM finetuned with CodePMP initialization achieves higher accuracy on both 1.5B and 7B models across mathematical and logical reasoning tasks, demonstrating that CodePMP enhances the model’s ability to differentiate correct from incorrect reasoning. Moreover, CodePMP exhibits strong generalization, yielding significant improvements across different reasoning tasks.
Table 1: Comparison of RM accuracies: reward models finetuned with CodePMP initialization achieve higher accuracies on the reasoning holdout test sets, demonstrating an improved ability to distinguish chosen responses from rejected ones.
Model PMP MathShepherd-pair Reclor-pair LogiQA2.0-pair
1.5B 0.7226 0.758 0.7538
0.8186 0.794 0.7774
7B 0.8777 0.862 0.8263
0.9274 0.874 0.8441
(a) BoN accuracies on mathematical reasoning.
(b) BoN (N=4) accuracies on logical reasoning.
Figure 3: Comparison of Best-of-N accuracies: reward models finetuned with CodePMP initialization consistently perform better, with improvements remaining robust as N increases, demonstrating CodePMP’s effectiveness in improving group-wise ranking capabilities.
(a) GSM8K / 1.5B
(b) MATH / 1.5B
(c) Reclor / 1.5B
(d) LogiQA2.0 / 1.5B
(e) GSM8K / 7B
(f) MATH / 7B
(g) Reclor / 7B
(h) LogiQA2.0 / 7B
Figure 4: Comparison of sample efficiency in RM finetuning: reward models finetuned with CodePMP initialization consistently achieve substantially higher Best-of-N accuracy when finetuning on the same amount of samples, demonstrating superior sample efficiency. Note that the horizontal axis grows exponentially with $\sqrt{2}$. The green lines represent the settings with CodePMP, while the blue lines represent the settings without CodePMP.
4.2.2 BoN Accuracy Results
We evaluate BoN accuracy across reasoning tasks with and without CodePMP initialization. As shown in Figure 8(a) and Table 8(b), RM finetuned with CodePMP initialization consistently achieves higher BoN accuracy across both mathematical and logical reasoning tasks for 1.5B and 7B models. This highlights CodePMP’s effectiveness in improving RM’s group-wise ranking performance.
Across different values of N, RM models initialized with CodePMP maintain their lead, showing robust improvement even as N increases to 256. In contrast, RM without CodePMP shows a sharp decline in accuracy as N increases, underscoring the stability CodePMP provides, likely due to the diverse code-preference pairs used during training.
For logical reasoning, the performance gap between CodePMP and non-PMP models is smaller, as $N$ is limited to 4, while in mathematical reasoning, $N$ reaches 256. This suggests that increasing $N$ further in logical reasoning could amplify the advantages of CodePMP in future evaluations.
4.2.3 RM Sample Efficiency Comparison
One key advantage of CodePMP is its ability to improve the sample efficiency of RM finetuning. To assess this, we conduct experiments with progressively larger sample sizes for RM finetuning. As indicated by (Kaplan et al., 2020), optimal results are achieved when the learning rate scheduler completes its decay at the end of training. Therefore, rather than evaluating intermediate checkpoints, we retrain models with varying sample sizes for optimal results. Figure 4 and  10 (Appendix E) show that as the sample size increases, RMs with CodePMP initialization consistently outperforms others in both BoN and RM accuracy. Notably, RMs finetuned with CodePMP initialization using just 0.5k samples surpasses RMs finetuned without CodePMP initialization using 40k samples on mathematical tasks, demonstrating CodePMP’s significant advantage in sample efficiency. However, as sample size increases, this advantage diminishes slightly, suggesting that with much larger datasets, CodePMP’s benefit may become less pronounced, but the cost of manual labeling remains a key consideration.
4.2.4 The Importance of Scalable PMP
A key benefit of using code data for PMP is the vast availability of publicly accessible, high-quality code-preference pairs, ensuring diversity. To validate scalability, we vary the number of training pairs for CodePMP and retrain models with different amounts of data. As shown in Figure 5, overall, increasing the number of code-preference pairs consistently improves BoN accuracy in both mathematical and logical reasoning tasks across model sizes, with no sign of diminishing returns. This indicates that further scaling the code-preference data would likely yield additional performance gains, underscoring the importance of building a scalable PMP pipeline.
(a) GSM8K / 1.5B
(b) MATH / 1.5B
(c) Reclor / 1.5B
(d) LogiQA2.0 / 1.5B
(e) GSM8K / 7B
(f) MATH / 7B
(g) Reclor / 7B
(h) LogiQA2.0 / 7B
Figure 5: Increasing the number of code-preference pairs consistently improves Best-of-N accuracy in both mathematical and logical reasoning tasks across model sizes, with no evident signs of diminishing returns. Note that the horizontal axis is scaled by $\sqrt{2}$, and the gray dashed line represents the results without CodePMP.
5 Ablation Studies
In this section, we present a detailed analysis of CodePMP design. Unless otherwise stated, all experiments used the 1B model due to resource limitations and present the results of mathematical reasoning due to page limitation. Results of logical reasoning refers to Appendix E.2.
5.1 Impact of Pair Construction
Model-Generated Pairs Comparison
We compare various pair construction methods generated by different models. In Figure 6(a), the samples before the “&” are positive, and those after are negative. “Source Code” refers to the original code snippet, while “1.3B-Des-Clip” indicates that 10% of the code description is removed before being input into a 1.3B CodeLLM to generate a rejected response. The green lines represent CodePMP’s choice. Results show that pairing positive samples from the 7B model with negative samples from the 1.5B model consistently delivers the best performance across all test sets. Given that code execution can generate reliable outputs, future work will explore incorporating execution feedback to create more accurate preference pairs.
Web-Crawled vs GitHub-Sourced Pairs
We also compare GitHub-sourced code with web-crawled data( Askell et al. (2021)) from platforms such as StackExchange and Reddit. As shown in Figure 7(a), GitHub-sourced pairs (“Source Code”) consistently outperform those from web platforms (“Webpage”), particularly as the number of solutions (N) increases. Moreover, the performance improvement of GitHub-sourced pairs shows no sign of plateauing, highlighting the importance of diverse, high-quality source code in building a scalable PMP pipeline.
(a) Different construction methods.
(b) Different pair sources.
Figure 6: Comparisons of BoN accuracy across different construction methods and pair sources.
(a) With or without EOC token.
(b) Using WSD or WCD learning rate schedulers.
Figure 7: Comparisons of BoN accuracy for different settings: EOC token and lr schedulers.
5.2 Impact of EOC Token
Experiments by( Askell et al. (2021)) show that adding an end-of-context (EOC) token to each sequence significantly improves overall performance. To explore its impact in the context of CodePMP, we compared performance with and without the EOC token. As shown in Figure 7(a), the EOC setting (“w/ EOC”) consistently underperform the setting without EOC tokens (“w/o EOC”) across different test tasks, which is opposed to ( Askell et al. (2021)) We attribute this discrepancy to the different model, data and evaluation settings.
5.3 Impact of Learning Rate Schedulers
In the CodePMP experiments, we use the warmup-stable-decay (WSD) learning rate scheduler( Hu et al. (2024)), which can effectively reduce the time required for scaling related experiments. Previous studies mainly employ a learning rate schedule with linear warmup followed by cosine decay, known as warmup-cosine decay (WCD). We compare the performance of WSD and WCD, as shown in Table 7(b), both schedulers yield similar results. Thus, to improve computational efficiency, we adopt the WSD scheduler for all experiments.
(a) BoN accuracies on mathematical reasoning.
(b) BoN (N=4) accuracies on logical reasoning.
Figure 8: CodePMP improves sample efficiency and Best-of-N performance of the Gemma-2B reward models on reasoning tasks, highlighting its broad applicability across diverse LLM architectures.
Table 2: Performance gains on coding RM and general RM (RMBench) evaluations show that CodePMP not only improves reasoning tasks but also generalizes well across various tasks.
RMBench
Model PMP Coding Summary Chat Chat Hard Safety Reasoning
1.5B 0.6841 0.4154 0.4804 0.5351 0.3665 0.2751
0.758 0.6126 0.9050 0.4364 0.3698 0.6041
7B 0.6912 0.5839 0.4972 0.5022 0.5240 0.6804
0.7619 0.7668 0.9413 0.5373 0.4906 0.9116
5.4 Validating CodePMP on Other LLMs
To further evaluate the generalizability of CodePMP, we validate its performance on the widely adopted Gemma-2B model (Team et al., 2024). As illustrated in Figure 8, the application of CodePMP results in significant performance gains in both mathemantical reasoning and logical reasoning evaluations. This not only underscores the robustness of CodePMP but also demonstrates its broad applicability in improving sample efficiency and overall performance across diverse LLM architectures.
5.5 Performance on Coding and General RM Benchmarks
We evaluate CodePMP on both code-specific and general reward modeling benchmarks. The CodeUltraFeedback_binarized test set serves as an in-domain evaluation, while RMBench provides an out-of-domain assessment. As shown in Table 2, models finetuned with CodePMP initialization consistently outperform those without CodePMP across various model sizes. These results demonstrate that CodePMP not only enhances performance in reasoning tasks but also generalizes well across a range of RM benchmarks.
6 Related Works
Reward Modeling
In the context of RLHF, reward models (RMs) have traditionally employed ranking models like Bradley-Terry and Plackett-Luce to represent human preferences (Bradley & Terry, 1952; Plackett, 1975; Cobbe et al., 2021b; Saunders et al., 2022; Lightman et al., 2023b; Wang et al., 2023b; Uesato et al., 2022; Luo et al., 2024; Yu et al., 2024; Stiennon et al., 2020; Nakano et al., 2021). More recently, probability-based approaches (Zhao et al., 2023; Jiang et al., 2023) have emerged, offering more precise predictions. Additionally, models such as Critique-out-Loud (Ankner et al., 2024) enhance RMs by integrating natural language feedback. Generative reward models (GRMs) (Yang et al., 2024b) further boost sample efficiency. Preference Modeling Pretraining (PMP) (Askell et al., 2021) introduces a novel pretraining phase, utilizing large-scale pairwise ranking data to enhance RM performance. Despite these advancements, many methods are hindered by the reliance on expensive manual annotations or limited datasets, constraining scalability. CodePMP mitigates this by automating preference data generation from code, significantly improving RM sample efficiency and reducing dependency on manual data collection.
Code Training
The inclusion of code in LLM pretraining has led to marked improvements in tasks such as commonsense reasoning (Madaan et al., 2022) and mathematical problem-solving (Liang et al., 2022; Shao et al., 2024; Yang et al., 2024a). Furthermore, code enhances general reasoning capabilities (Muennighoff et al., 2023; Fu & Khot, 2022; Ma et al., 2023). Recent studies (Dong et al., 2023; Ma et al., 2023) indicate that incorporating code during supervised finetuning strengthens LLMs, particularly in complex decision-making tasks. CodePMP takes a pioneering approach by utilizing scalable, synthetically generated code preference pairs, reducing the dependence on manual annotation (Dubey et al., 2024; Gemini-Team et al., 2024; Groeneveld et al., 2024; Bi et al., 2024). This methodology enhances sample efficiency and scalability in reasoning-intensive tasks, presenting new opportunities for further improving LLM performance.
LLM Reasoning
Improving reasoning capabilities in LLMs remains a significant challenge, with various advanced methods being proposed. Chain of Thought (CoT) prompting (Wei et al., 2022; Fu et al., 2023) improves reasoning by generating intermediate steps, while CoT combined with supervised finetuning (SFT) further enhances performance (Cobbe et al., 2021a; Liu et al., 2024; Yu et al., 2023). Other approaches focus on expanding inference time computation, such as problem decomposition (Zhou et al., 2022), search-based methods like MCTS (Xu, 2023), and using LLMs as verifiers (Huang et al., 2022; Luo et al., 2023). Reward models, including outcome-based (ORM) and process-based (PRM), have also shown success, with PRM delivering superior results (Lightman et al., 2023a; Wang et al., 2023b). Distinct from these approaches, CodePMP introduces a scalable preference model pretraining phase that can integrate seamlessly with all the aforementioned techniques.
7 Conclusion & Future Works
This paper introduces CodePMP, a scalable pretraining approach that leverages code-preference pairs to improve reasoning capabilities in large language models. Experimental results validate that CodePMP significantly improves sample efficiency and boosts performance on reasoning tasks.
For future work, we aim to extend CodePMP in two key directions. CodePrMP will focus on utilizing compiler and interpreter verifiability to provide low-cost process supervision signals. GenPMP will explore how to improve sample efficiency and the performance of generative reward models by integrating code data.
Appendix A Hyperparameters
We outline key hyperparameters in this Table 3. In the tables, WSD refers to the warmup-stable-decay learning rate scheduler( (Hu et al., 2024)), which has the benefit of reducing the time required for scaling law experiments.
Table 3: Hyperparameters for CodePMP training, mathematical reasoning RM finetuning, and logical reasoning RM finetuning.
CodePMP Mathematical RM Logical RM
HP 1.5B 7B 1.5B 7B 1.5B 7B
epoch 1 1 1 1 1 1
bs 1024 1024 64 64 64 64
lr 3e-6 1e-6 1e-6 3e-7 1e-5 1e-5
lr scheduler WSD WSD WCD WCD WCD WCD
warmup ratio 0.03 0.03 0.03 0.03 0.25 0.25
decay ratio 0.1 0.1 - - - -
weight decay 0.1 0.1 0 0 0 0
max length 1024 1024 1024 1024 1024 1024
Appendix B Training Pipeline
Figure B presents an overview of the complete training pipeline. The process begins with base language model (LM) pretraining on trillions of tokens from general text, followed by a preference model pretraining (PMP) phase using billions of tokens from code preference pairs. Finally, the model is finetuned on a smaller, more specialized dataset relevant to reasoning tasks, typically consisting of millions of tokens.
Figure 9: An overview of the complete training pipeline.
Appendix C CodePMP Dataset
As shown in Table 4, the constructed CodePMP dataset consists of a total of 28 million files, accounting for 19 billion tokens across various languages. The dataset is primarily composed of Python files, with 20 million files and 13.1 billion tokens, followed by Notebook files, contributing 3 million files and 2.1 billion tokens, and other programming languages with 5 million files and 3.8 billion tokens. This diverse dataset supports the pretraining phase, aiding the model in generalizing across multiple reasoning tasks.
The average length of chosen responses varies slightly between different languages. Python files exhibit an average response length of 170 tokens for chosen responses and 167 tokens for rejected responses, while Notebook files have slightly shorter average lengths of 158 and 155.5 tokens, respectively. Other languages show the highest average response lengths, with 213.2 tokens for chosen and 210 tokens for rejected responses. Overall, the average length of chosen responses across the dataset is 194.5 tokens, while the average length of rejected responses is 189.9 tokens, indicating minimal bias in the dataset based on response length.
Table 4: Amount and average length statistics of CodePMP dataset.
Amount Statistics Average Length
Language Files (M) Tokens (B) Chosen Rejected
Python 20 13.1 170.0 167.0
Notebook 3 2.1 158.0 155.5
Other Languages 5 3.8 213.2 210.0
Total 28 19.0 194.5 189.9
Appendix D RM Finetuning Dataset
D.1 Mathematical Reasoning
The RM finetuning for mathematical reasoning uses the MathShepherd dataset( (Wang et al., 2023b)), which contains 444k query-response samples, with some queries having multiple distinct responses. We divide the dataset into a 400k training set and a 44k test set. For RM finetuning, we construct preference pairs by selecting both correct and incorrect responses for the same query. To form the 4.3k test set, we combine one positive and negative sample for each query from the original test set.
We also create two training sets of different sizes: MathShepherd-preference-800k and MathShepherd-preference-40k. The 800k training set is built by combining multiple positive and negative samples for each query in the original training set, resulting in 800k samples. In contrast, the 40k training set randomly selects one positive-negative pair for each query, totaling 40k samples.
D.2 Logical Reasoning
D.2.1 Reclor
Reclor is a human-annotated reading comprehension reasoning dataset, where each sample consists of a passage, a question, and multiple options. To create preference pairs, we combine the correct and incorrect options for the same question. This process results in a total of 14.5k preference pairs, with 14k pairs used for training and 1.5k for testing, forming the Reclor-preference dataset.
D.2.2 LogiQA2.0
For logical reasoning finetuning scaling analysis, we synthesize a preference dataset based on the logical reasoning dataset LogiQA2.0. LogiQA2.0 is a reading comprehension benchmark requiring discrete reasoning over passages, with 96k crowdsourced, adversarially created questions. To answer correctly, models must resolve references (which may point to multiple locations in the input) and perform discrete operations like addition, counting, or sorting. We use four models (Qwen2-7B-Instruct [^1] https://huggingface.co/Qwen/Qwen2-7B-Instruct, Qwen2-72B-Instruct [^2] https://huggingface.co/Qwen/Qwen2-72B-Instruct, DeepSeek-V2-Chat [^3] https://huggingface.co/deepseek-ai/DeepSeek-V2-Chat, and DeepSeek-Coder-V2-Instruct [^4] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Instruct) to sample queries from the LogiQA2.0 dataset multiple times, with sampling $topp=1$, and $topk\in\{0,0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$. Answer correctness is annotated using DeepSeek-7B-Math-Compare-Answer [^5] https://huggingface.co/Tianqiao/DeepSeek-7B-Math-Compare-Answer. Correct and incorrect answers are combined to create preference pairs, resulting in 1,019k pairs, with 977k used for training and 42k for testing, forming the LogiQA2.0 preference dataset for RM scaling analysis.
D.3 CodeUltraFeedback_binarized
CodeUltraFeedback_binarized [^6] https://huggingface.co/datasets/coseal/CodeUltraFeedback_binarized is a preference dataset in the code domain, consisting of 9.5k preference pairs. We randomly split the dataset, using 90% of the samples for finetuning training and 10% for testing RM accuracy.
Appendix E More Experimental Results
Due to the length of the paper, we only report important experimental results in the experiment and ablation sections. In this section, we present complementary results in more detail.
E.1 RM Sample Efficiency Comparison: RM accuracy results
(a) MathShepherd-pair / 1.5B
(b) Reclor-pair / 1.5B
(c) LogiQA2.0-pair / 1.5B
(d) MathShepherd-pair / 7B
(e) Reclor-pair / 7B
(f) LogiQA2.0-pair / 7B
Figure 10: Comparison of sample efficiency of RM finetuning: Trends of RM accuracy with sample size increases. Note: The horizontal axis increases exponentially with $\sqrt{2}$. Therefore, the further the data intervals, the more RM data is effectively saved. We use different colors to highlight the results of different model sizes.
We use RM accuracy evaluation to compare the sample efficiency of RM finetuning, as shown in Figure 10. Consistent with the conclusion from the main experiment, CodePMP improves sample efficiency. Under the same sample conditions, it consistently delivers stable improvements, making RM training more efficient.
E.2 Ablation Results on logical Reasoning
Table 5 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for different Model-generated pair construction methods for CodePMP. Overall, the $6.7B\&1.3B$ setting is the best among them.
Table 6 shows the comparison of the BoN accuracy performance of ReClor and LogiQA2.0 using GitHub source code and web crawled code data to build CodePMP training pairs. The results verify that Github code is a better source for pair construction.
Table 7 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for CodePMP models with and without EOC. In general, the setup without using EOC brings better results.
Table 8 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for WCD and WSD lr scheduler settings. The results below show that WSD lr scheduler brings better results than WCD.
Table 5: Comparison of ReClor and LogiQA2.0 BoN accuracies for different Model-generated pair construction methods for CodePMP. Overall, the $6.7B\&1.3B$ setting is the best among them.
Test Set
RM Finetune Constructions Method ReClor LogiQA2.0
ReClor-pair Source Code & 1.3B 0.550 0.4464
Source Code & 1.3B-Des-Clip 0.600 0.4522
Source Code & 6.7B 0.578 0.4605
1.3B & 1.3B-Des-Clip 0.572 0.4745
6.7B & 1.3B-Des-Clip 0.564 0.4809
6.7B & 1.3B 0.608 0.5032
LogiQA2.0-pair Source Code & 1.3B 0.708 0.5217
Source Code & 1.3B-Des-Clip 0.704 0.5268
Source Code & 6.7B 0.714 0.5198
1.3B & 1.3B-Des-Clip 0.738 0.5402
6.7B & 1.3B-Des-Clip 0.742 0.5274
6.7B & 1.3B 0.734 0.5415
Table 6: Comparison of the BoN accuracy performance of ReClor and LogiQA2.0 using GitHub source code and web crawled code data to build CodePMP training pairs. The results below show that Github code is a better source for pair construction.
Test Set
RM Finetune Data Source ReClor LogiQA2.0
ReClor-pair Webpage 0.574 0.4898
Github 0.582 0.4981
LogiQA2.0-pair Webpage 0.742 0.5293
Github 0.752 0.5504
Table 7: Comparison of ReClor and LogiQA2.0 BoN accuracies for CodePMP models with and without EOC. In general, the setup without using EOC brings better results.
Test Set
RM Finetune Model ReClor LogiQA2.0
ReClor-pair w/o EOC 0.582 0.4981
w/ EOC 0.596 0.4617
LogiQA2.0-pair w/o EOC 0.752 0.5504
w/ EOC 0.686 0.4809
Table 8: Comparison of ReClor and LogiQA2.0 BoN accuracies for WCD and WSD lr scheduler settings. The results below show that WSD lr scheduler brings better results than WCD.
Test Set
RM Finetune Model ReClor LogiQA2.0
ReClor-pair WCD 0.552 0.4828
WSD 0.582 0.4981
LogiQA2.0-pair WCD 0.748 0.5204
WSD 0.752 0.5504
Table 9: Cross-validation evaluation: ReClor and LogiQA2.0 BoN accuracies, black numbers are the results of cross-validation. In the cross-validation evaluation, CodePMP can still bring stable improvements on different test sets.
Test Set
RM Finetune Model PMP ReClor LogiQA2.0
ReClor-pair 1.5B 0.534 0.4592
0.582 0.4981
7B 0.724 0.5453
0.756 0.5835
LogiQA2.0-pair 1.5B 0.710 0.5293
0.752 0.5504
7B 0.748 0.6371
0.794 0.6779
E.3 Cross-validations on logical Reasoning
We further conduct cross-dataset validation on logical reasoning. Specifically, we evaluate the RM models finetuned on the ReClor-pairs dataset using the LogiQA2.0 test set. Similarly, we evaluate the RM models finetuned on the LogiQA2.0-pairs dataset using the ReClor test set. As shown in Table 9, CodePMP consistently improves RM evaluation performance, demonstrating that the enhancements CodePMP brings to RM training are robust and generalizable. Note that the BoN accuracies of the RM trained with LogiQA2.0-pair on the Reclor test set are higher than those of the RM trained directly on the Reclor-pair, because the LogiQA2.0-pair dataset is three times larger than the Reclor-pair dataset.
Appendix F Logical Reasoning Evaluation Examples
We randomly select and present examples from the Reclor test set, which consists of multiple-choice questions based on a given passage. While it is possible to have the model generate additional candidate answers to create a Best-of-N test, it becomes difficult to ensure that the original correct answer remains among the options after introducing new candidates, and to identify the new correct answer. We attempt to use GPT-4o to annotate the correct answers for 32 responses, but the consistency with manual inspection is low, as is the consistency of GPT-4o’s own multiple annotations. It can be inferred that the consistency rate would worsen if expanded to 256 responses. Therefore, after careful consideration, we decide to use RM to score only the original four manually annotated answer options, match the top-ranked option with the manually annotated correct answer, and calculate accuracy. In principle, this method is equivalent to the Best-of-4 test.
Table 10: Examples from the Reclor test set, which consists of multiple-choice questions based on a given passage.
ID Text Options Answer
12824 Mayor: Four years ago, when we reorganized the city police department in order to save money, critics claimed that the reorganization would make the police less responsive to citizens and would thus lead to more crime. The police have compiled theft statistics from the years following the reorganization that show that the critics were wrong. There was an overall decrease in reports of thefts of all kinds, including small thefts.
Question: Which of the following, if true, most seriously challenges the mayor’s argument?
1. In other cities where police departments have been similarly reorganized, the numbers of reported thefts have generally risen following reorganization. 2. When city police are perceived as unresponsive, victims of theft are less likely to report thefts to the police. 3. The mayor’s critics generally agree that police statistics concerning crime reports provide the most reliable available data on crime rates. 4. The mayor’s reorganization of the police department failed to save as much money as it was intended to save. 1
218 Jupiter is a gas giant planet and the largest planet in the solar system. Its mass is 2.5 times the total mass of the other seven planets in the solar system. Observations have found that most of the more than 70 moons surrounding Jupiter are composed of water ice. Therefore, Jupiter’s atmosphere should contain a considerable amount of water.
Question: Which of the followings, if true, can best support the above statement?
1. After hundreds of millions of years, the satellite may slowly fall onto the planet. 2. Many of the water in interstellar space exists in gaseous form. 3. Uranus is also a gas giant planet, and it has been confirmed that it contains a lot of water ice. 4. The satellite and the planets around it were formed from the same gas and dust at the same time. 3
10376 Lake Dali is a barrier lake on the plateau formed by volcanic eruptions. Like salmon living in the sea, Hua Zi fish-Leuciscus waleckii, which lives in a brackish lake, must migrate to the upper reaches of the Tanshui River to spawn and breed, although the four rivers currently flowing into Lake Dali are inland rivers, and none of them leads to the sea. Scientists are still convinced that the Huaziyu in Lake Dali first migrated from the ocean.
Question: Which of the following options, if true, provides the best explanation for scientists’ beliefs?
1. The Leuciscus waleckii that lives in the waters such as Heilongjiang is twice as big as the Leuciscus waleckii fish in Lake Dari. 2. The caught Hua Zi fish can only survive for a day or two after being put into sea water or fresh water, and will decay quickly after death. 3. Melting glaciers will form Lake Dali, and the overflowing lake was once connected to the Liao River, which flowed into the ocean. 4. The researchers put the fry of Hua Zi fish in Dali Lake into Gainao thousands of miles away, and the culture was successful. 2
13334 It is repeatedly claimed that the dumping of nuclear waste poses no threat to people living nearby. If this claim could be made with certainty, there would be no reason for not locating sites in areas of dense population. But the policy of dumping nuclear waste only in the more sparsely populated regions indicates, at the very least, some misgiving about safety on the part of those responsible for policy.
Question: Which one of the following, if true, would most seriously weaken the argument?
1. Until there is no shred of doubt that nuclear dumps are safe, it makes sense to situate them where they pose the least threat to the public. 2. There are dangers associated with chemical waste, and it, too, is dumped away from areas of dense population. 3. In the event of an accident, it is certain that fewer people would be harmed in a sparsely populated than in a densely populated area. 4. Dumping of nuclear waste poses fewer economic and bureaucratic problems in sparsely populated than in densely populated areas. 3

### v2

@@
 We evaluate CodePMP on mathematical reasoning tasks (GSM8K, MATH) and logical reasoning tasks (ReClor, LogiQA2.0), consistently showing significant improvements in reasoning performance of LLMs and highlighting the importance of scalable preference model pretraining for efficient reward modeling.
+CodePMP: Scalable Preference Model Pretraining for Large Language Model Reasoning
+Huimu Yu1,2, Xing Wuthanks: Equal contribution.
+This work is jointly completed by Huimu Yu under the guidance of Xing Wu.1,2,3 🖂, Haotian Xu3, Debing Zhang3, Songlin Hu1,2 🖂 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Xiaohongshu Inc {yuhunmu,wuxing,husonglin}@iie.ac.cn, {xuhaotian,dengyang}@xiaohongshu.com
 1 Introduction
+Large language models (LLMs) have achieved remarkable progress in natural language understanding and generation, driven by advancements in scalable pretraining and finetuning techniques, including supervised finetuning (SFT) (Wang et al., 2022, 2023a) and Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022a; Lightman et al., 2023; Bai et al., 2022b; Gulcehre et al., 2023; Schulman et al., 2017; Rafailov et al., 2024).
+Despite these advances, enhancing LLMs’ reasoning capabilities, particularly for complex logical and mathematical tasks, remains a significant challenge (Wang et al., 2023b; Zhang et al., 2024b).
+While RLHF has proven effective for improving model performance, its efficacy is constrained by the availability of high-quality preference data, which is expensive and labor-intensive to collect (Cobbe et al., 2021; Zheng et al., 2024).
+This limitation impedes the scalability of reward model (RM) finetuning, which is instrumental in guiding LLMs toward optimal outputs.
 Figure 1: Compared to directly finetuning reward models, CodePMP significantly improves the sample efficiency and capability of reward models, which in turn boosts the generator’s(MetaMath-Mistral-7B) reasoning performance (Best-of-N accuracy) across both mathematical reasoning tasks (GSM8K and MATH) and logical reasoning tasks (ReClor and LogiQA2.0).
-Large language models (LLMs) have made remarkable progress in natural language understanding and generation, benefiting from scalable pretraining and finetuning techniques like supervised finetuning (SFT) (Wang et al., 2022; 2023a) and Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022a; Lightman et al., 2023b; Bai et al., 2022b; Gulcehre et al., 2023; Schulman et al., 2017; Rafailov et al., 2024).
-However, enhancing LLMs’ reasoning abilities, particularly in complex logical and mathematical tasks, remains a significant challenge (Wang et al., 2023b; Zhang et al., 2024b).
-Although RLHF is effective, it relies heavily on high-quality preference data, which is costly and labor-intensive to annotate (Cobbe et al., 2021b; Zheng et al., 2024).
-This limitation impedes the scalability of reward model (RM) finetuning, which is essential for guiding LLMs toward optimal outputs.
 To alleviate this issue, prior works like Anthropic’s Preference Model Pretraining (PMP) (Askell et al., 2021) have proposed improving reward modeling data efficiency by pretraining preference models on large-scale preference data from public sources like Reddit and Wikipedia, followed by an efficient finetuning on limited high-quality human-annotated data.
 However, this approach is less effective for reasoning tasks due to the scarcity of reasoning preference pairs available online.
-Compared to other tasks, manually annotating preference data for reasoning is inherently more challenging and difficult to scale (Zhang et al., 2024b; Zhou et al., 2023), highlighting the urgent need for a scalable PMP approach for reasoning tasks.
-Figure 2: An illustration of the CodePMP process.
-First, raw code collected from GitHub is cleaned and summarized into code prompts (descriptions).
-Then, for each code prompt, a weak CodeLLM generates a rejected response, while a stronger CodeLLM produces a chosen response.
-Finally, these <chosen, rejected> pairs, accumulated in the millions, form the pretraining dataset for the preference model.
-This pretraining process improves not only sample efficiency but also the performance for downstream reasoning reward model finetuning.
+Compared to other tasks, manually annotating preference data for reasoning is inherently more challenging to scale (Zhang et al., 2024b; Zhou et al., 2023), highlighting the urgent need for a scalable PMP approach for reasoning tasks.
+[…]
 In this paper, we propose CodePMP, a scalable preference model pretraining pipeline that enhances LLM reasoning abilities using synthesized preference pairs derived from high-quality, publicly available source code.
@@
 Recent works (Zhang et al., 2024b; Aryabumi et al., 2024) also show a strong correlation between code training and reasoning improvements in LLMs.
-By leveraging the huge amount and diverse coverage of source code available on platforms like GitHub, CodePMP offers a scalable solution for pretraining preference models, thereby improving RM finetuning efficiency and enhancing LLMs reasoning performance.
-Specifically, CodePMP generates preference pairs by synthesizing chosen and rejected code responses for a given code-related prompt or description using CodeLLMs.
-A strong CodeLLM produces higher-quality (chosen) responses, while a weaker model generates suboptimal or even low-quality (rejected) responses.
-These <chosen, rejected> pairs, accumulated in the millions, form a large-scale synthesized preference dataset.
-This dataset is then used to pretrain the preference model with pairwise ranking objectives (Cobbe et al., 2021b; Charniak & Johnson, 2005), providing an good initialization for further fine-tuning the reward models.
-We evaluate CodePMP on widely studied reasoning tasks, including mathematical reasoning tasks such as GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021), as well as logical reasoning tasks like ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023).
-Our experiments demonstrate that CodePMP significantly improves RM fine-tuning accuracy and Best-of-N performance in reasoning tasks, outperforming direct RM fine-tuning, as highlighted in Figure 1.
-Moreover, additional experimental results reveals that RMs initialized with CodePMP exhibit greater robustness across different tasks.
+By leveraging the huge amount and diverse coverage of source code available on platforms like GitHub, CodePMP offers a scalable solution for pretraining preference models, thereby improving RM finetuning efficiency and enhancing LLMs’ reasoning performance.
+[…]
+We evaluate CodePMP on widely studied reasoning tasks, including mathematical reasoning tasks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), as well as logical reasoning tasks like ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023).
+Our experiments show that CodePMP significantly improves RM finetuning accuracy and Best-of-N performance in reasoning tasks, outperforming direct RM finetuning, as highlighted in Figure 1.
+Moreover, additional results reveal that RMs initialized with CodePMP exhibit greater robustness across different tasks.
 These results indicate that code-derived preference data provides a scalable, cost-effective solution for enhancing LLM reasoning capabilities while reducing reliance on extensive preference annotation, achieving more effective reward modeling for reasoning tasks.
@@
 Language Modeling
-Language modeling (LM) is a fundamental task in natural language processing, aimed at modeling sequences of language.
-This is typically achieved through Causal Language Models (Causal LM), where the model is trained to maximize the likelihood of predicting the next word $w_{t}$ given the preceding words $w_{1},w_{2},\dots,w_{t-1}$.
-The training process minimizes the negative log-likelihood of the predicted word sequence:
-$\mathcal{L}_{\text{LM}}=-\sum_{t=1}^{T}\log P(w_{t}|w_{1},w_{2},\dots,w_{t-1})$
-This loss function $\mathcal{L}_{\text{LM}}$ encourages the model to capture the underlying patterns in the data.
-Transformer architectures (Vaswani, 2017) have become the standard for Causal LM due to their ability to handle long-range dependencies effectively.
+Language modeling represents a fundamental task in natural language processing aimed at modeling sequential language data.
+This is typically implemented through Causal Language Models (Causal LM), which maximize the likelihood of predicting the next token $w_{t}$ given preceding tokens $w_{1},w_{2},\dots,w_{t-1}$.
+The training process minimizes the negative log-likelihood:
+$\mathcal{L}_{\text{LM}}=-\sum_{t=1}^{T}\log P(w_{t}|w_{1},w_{2},\dots,w_{t-1})$ (1)
+This loss function $\mathcal{L}_{\text{LM}}$ encourages the model to capture underlying patterns in the data.
+Transformer architectures (Vaswani, 2017) are the standard for Causal LM due to their ability to handle long-range dependencies effectively.
 Reward Modeling
-Reward modeling (RM) is crucial in reinforcement learning from human feedback (RLHF), providing a scalar reward signal that guides the learning process based on output quality.
-The reward model $R_{\theta}$ predicts the quality of an output $y$ given a context $x$ as $s=R_{\theta}(x,y)$, where $s$ is the scalar reward score.
-In preference modeling, RM predicts the relative quality of outputs by comparing pairs.
-A common method is the Pairwise Ranking Loss, where the model assigns higher scores to preferred (chosen) outputs:
-$\mathcal{L}_{\text{RM}}=-\log\left(\sigma(s_{\text{chosen}}-s_{\text{rejected}})\right)$
+Reward modeling (RM) is integral to reinforcement learning from human feedback (RLHF), providing scalar reward signals that guide learning based on output quality.
+The reward model $R_{\theta}$ predicts the quality of an output $y$ given a context $x$ as $s=R_{\theta}(x,y)$.
+In preference modeling, RMs predict relative quality by comparing output pairs.
+A standard approach employs the Pairwise Ranking Loss, which assigns higher scores to preferred (chosen) outputs:
+$\mathcal{L}_{\text{RM}}=-\log\left(\sigma(s_{\text{chosen}}-s_{\text{rejected}})\right)$ (2)
 , where $s_{\text{chosen}}=R_{\theta}(x,y_{\text{chosen}})$ and $s_{\text{rejected}}=R_{\theta}(x,y_{\text{rejected}})$, and $\sigma(\cdot)$ is the sigmoid function.
 Best-of-N Sampling
-Best-of-N (BoN) sampling improves LLM reasoning (Cobbe et al., 2021b; Lightman et al., 2023b).
-In this approach, $N$ candidate solutions $\{y_{1},y_{2},\dots,y_{N}\}$ are generated by sampling from the LLM’s output distribution for a given problem.
-A reward model scores each candidate and selects the highest-scoring one as the final answer:
-$\hat{y}=\arg\max_{y_{i}\in\{y_{1},y_{2},\dots,y_{N}\}}R_{\theta}(x,y_{i})$
+Best-of-N (BoN) sampling enhances LLM reasoning (Cobbe et al., 2021; Lightman et al., 2023) by generating $N$ candidate solutions $\{y_{1},y_{2},\dots,y_{N}\}$ for a given problem, then using a reward model to score and select the highest-scoring candidate:
+$\hat{y}=\arg\max_{y_{i}\in\{y_{1},y_{2},\dots,y_{N}\}}R_{\theta}(x,y_{i})$ (3)
 , where $R_{\theta}(x,y_{i})$ represents the reward score for each candidate $y_{i}$.
@@
 3.1 Model Design
-Code Preference Model Pretraining (CodePMP) is designed to enhance the sample efficiency of reward models, particularly for reasoning tasks where high-quality preference data is scarce.
-Traditionally, reward models are finetuned on small, curated datasets, which limits their effectiveness in complex tasks like mathematical reasoning or logical deduction.
+Code Preference Model Pretraining (CodePMP) enhances the sample efficiency of reward models, particularly for reasoning tasks where high-quality preference data is scarce.
+Traditionally, reward models are finetuned on small, curated datasets, limiting their effectiveness in complex tasks like mathematical reasoning or logical deduction.
 CodePMP mitigates this limitation by introducing a pretraining phase between basic language model pretraining and finetuning on domain-specific reasoning datasets.
 This phase leverages a large, diverse dataset of code-preference pairs, enabling the model to learn generalizable patterns and ranking strategies.
-CodePMP training involves two key components: Reward Modeling (RM) and Language Modeling (LM).
+CodePMP training involves two components: Reward Modeling (RM) and Language Modeling (LM).
 In RM, the model is trained on code-preference pairs, learning to assign higher scores to the chosen code through a pairwise ranking loss.
 In LM, only the chosen code is used for autoregressive training to maintain the model’s general capabilities.
-The overall loss is a combination of the RM and LM losses, ensuring the model enhances its ranking ability without sacrificing general language modeling performance: $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{LM}}$.
+The overall loss combines the RM and LM losses, ensuring the model enhances its ranking ability without sacrificing general language modeling performance: $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{RM}}+\mathcal{L}_{\text{LM}}$.
 3.2 Data Construction
-To enable scalable preference model pretraining, we construct a dataset sourced from GitHub, which includes a diverse range of repositories and associated metadata.
-The dataset consists of two primary components: Repository Data comprises over 1.3 billion code files from GitHub repositories, while GitHub Metadata includes information such as commit histories, discussions, pull requests, and issues.
+To enable scalable preference model pretraining, we construct a dataset sourced from GitHub, containing over 1.3 billion code files from GitHub repositories.
 The CodePMP dataset is constructed through a systematic process.
-First, raw source code is processed by a description summarizer, typically an instruction-tuned CodeLLM, to generate prompts that describe the functionality of the code.
-These prompts are then used by two CodeLLMs of different capabilities to generate code snippets:
+First, raw source code is processed by a description summarizer—typically an instruction-tuned CodeLLM—to generate prompts describing the code’s functionality.
+Two CodeLLMs with different capabilities then generate code snippets based on these prompts:
   • Chosen response: Generated by a more advanced CodeLLM (e.g., 6.7B parameters).
@@
 This scalable approach significantly enhances pretraining efficiency, improving performance on downstream tasks.
-The steps of the CodePMP methodology are outlined systematically in Algorithm 1.
-Algorithm 1 Code Preference Model Pretraining (CodePMP)
-1:Source code repository $S$, Strong CodeLLM $M_{\text{strong}}$, Weak CodeLLM $M_{\text{weak}}$
-2:Pretrained Model
-3:Input: Source code $S$
-4:Summarize description $D$ using $M_{\text{strong}}$ on $S$
-5:for each $D_{i}\in D$ do
-6:     Generate Chosen Response using $M_{\text{strong}}$
-7:     Generate Rejected Response using $M_{\text{weak}}$
-8:end for
-9:Calculate LM Loss $\mathcal{L}_{\text{LM}}$ on Response
-10:Calculate RM Loss $\mathcal{L}_{\text{rank}}$ using Chosen Response and Rejected Response
-11:Train PMP Model using $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{rank}}+\mathcal{L}_{\text{LM}}$
-4 Experimental
-In this section, we first outline the experimental setup, followed by the experimental results, highlighting that CodePMP is a highly scalable method.
+The steps of the CodePMP methodology are outlined systematically in Figure 2.
+4 Experiments
+In this section, we outline the experimental setup and then the experimental results, highlighting that CodePMP is a highly scalable method.
+PMP
+MathShepherd
+-pair
+Reclor
+-pair
+LogiQA2.0
+-pair
+Qwen2-1.5B
+0.7226 0.758 0.7538
+0.8186 0.794 0.7774
+Qwen2-7B
+0.8777 0.862 0.8263
+0.9274 0.874 0.8441
+Table 1: Reward model accuracy comparison: CodePMP-initialized models perform better on reasoning test sets, showing better discrimination ability.
 4.1 Experimental Settings
@@
 Data Construction
-We generate code preference pairs following Algorithm 1, using the deepseek-coder-6.7b-instruct model as the strong CodeLLM to generate chosen responses and the deepseek-coder-1.3b-instruct model as the weak CodeLLM to generate rejected responses.
+We generate code preference pairs by using the deepseek-coder-6.7b-instruct model as the strong CodeLLM to generate chosen responses and the deepseek-coder-1.3b-instruct model as the weak CodeLLM to generate rejected responses.
 The constructed CodePMP dataset includes 28 million files and 19 billion tokens.
@@
 In addition, the average lengths of the chosen and rejected responses are similar, ensuring that response length does not bias the CodePMP learning process.
-Details are provided in Appendix C and Table 4.
+Details are provided in Appendix.
 CodePMP Training
-By default, we initialize the preference models with the publicly available Qwen models (Yang et al., 2024a), using different model sizes, specifically Qwen2-1.5B and Qwen2-7B.
-Detailed hyperparameters for CodePMP training are provided in Appendix B.
+By default, we initialize the preference models with the publicly available Qwen models (Yang et al., 2024), using different model sizes, specifically Qwen2-1.5B and Qwen2-7B.
+Detailed hyperparameters for CodePMP training are provided in Appendix.
 4.1.2 Reasoning Finetuning Settings
-We validate CodePMP on reward models across two reasoning task types: mathematical and logical reasoning.
-The reward model is finetuned on corresponding preference datasets for each task.
-For mathematical reasoning, we use the MathShepherd-pair dataset, derived from MathShepherd (Wang et al., 2023b), and evaluate the model on a holdout test set to assess RM accuracy.
-Similarly, for logical reasoning, we use the ReClor-pair and LogiQA2.0-pair datasets, derived from ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023), respectively.
-We train reward models on these datasets, with holdout test sets used to evaluate model accuracy.
-Dataset construction and finetuning hyperparameters are provided in Appendix D and  B.
+We evaluate CodePMP on mathematical and logical reasoning tasks using dedicated preference datasets.
+For mathematical reasoning, we finetune reward models on MathShepherd-pair dataset, derived from MathShepherd (Wang et al., 2023b), while logical reasoning models use ReClor-pair and LogiQA2.0-pair datasets, derived from ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023) respectively.
+Each model is finetuned on its corresponding training set and evaluated on its respective holdout test set for accuracy assessment.
+Implementation details for dataset construction and hyperparameters are provided in Appendix.
 4.1.3 Evaluation Settings
-Following (Zhang et al., 2024a), we adopt two evaluation metrics:
-RM Accuracy
-This metric measures the accuracy of the reward model in distinguishing chosen from rejected solutions on the holdout test sets.
-It provides insight into the model’s ability to classify individual sequences.
-Best-of-N (BoN) Accuracy
-This metric evaluates the proportion of correct solutions selected by the finetuned RM from $N$ candidate responses.
-It assesses the model’s group-wise ranking performance, focusing on its ability to select the correct answer from a set of candidates.
-We use MetaMath-Mistral-7B (Yu et al., 2023) as the generator for this evaluation.
-For mathematical reasoning, we use the GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021) test sets.
-For logical reasoning, we evaluate on the ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023) test sets.
-Further details can be found in Appendix D.
-Note that logical reasoning questions typically involve a paragraph followed by statements to be judged true or false, making Best-of-N evaluation challenging.
-Therefore, we use multiple-choice accuracy, where the reward model ranks four manually annotated options and selects the best one.
-This metric is equivalent to Best-of-4, and thus, for logical reasoning tasks, multiple-choice accuracy and Best-of-N are used interchangeably.
+Following (Zhang et al., 2024a), we evaluate using two metrics: (1) RM Accuracy measures the reward model’s ability to distinguish chosen from rejected solutions on holdout test sets, providing insight into the model’s ability to classify individual sequences; and (2) Best-of-N (BoN) Accuracy assesses the percentage of correct solutions selected by the RM from $N$ candidate responses, evaluating the model’s group-wise ranking performance and ability to identify the best answer from multiple candidates.
+We use MetaMath-Mistral-7B (Yu et al., 2023) as the generator for BoN evaluation.
+We evaluate on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) for mathematical reasoning, and ReClor (Yu et al., 2020) and LogiQA2.0 (Liu et al., 2023) for logical reasoning.
+For logical reasoning tasks, we use multiple-choice accuracy (equivalent to Best-of-4) where the RM ranks four manually annotated options, as logical reasoning questions typically consist of paragraphs followed by statements to be judged, making standard BoN evaluation challenging.
 4.2 Experimental Results
@@
 Moreover, CodePMP exhibits strong generalization, yielding significant improvements across different reasoning tasks.
-Table 1: Comparison of RM accuracies: reward models finetuned with CodePMP initialization achieve higher accuracies on the reasoning holdout test sets, demonstrating an improved ability to distinguish chosen responses from rejected ones.
-Model PMP MathShepherd-pair Reclor-pair LogiQA2.0-pair
-1.5B 0.7226 0.758 0.7538
-0.8186 0.794 0.7774
-7B 0.8777 0.862 0.8263
-0.9274 0.874 0.8441
 (a) BoN accuracies on mathematical reasoning.
 (b) BoN (N=4) accuracies on logical reasoning.
-Figure 3: Comparison of Best-of-N accuracies: reward models finetuned with CodePMP initialization consistently perform better, with improvements remaining robust as N increases, demonstrating CodePMP’s effectiveness in improving group-wise ranking capabilities.
-(a) GSM8K / 1.5B
-(b) MATH / 1.5B
-(c) Reclor / 1.5B
-(d) LogiQA2.0 / 1.5B
-(e) GSM8K / 7B
-(f) MATH / 7B
-(g) Reclor / 7B
-(h) LogiQA2.0 / 7B
-Figure 4: Comparison of sample efficiency in RM finetuning: reward models finetuned with CodePMP initialization consistently achieve substantially higher Best-of-N accuracy when finetuning on the same amount of samples, demonstrating superior sample efficiency.
-Note that the horizontal axis grows exponentially with $\sqrt{2}$.
-The green lines represent the settings with CodePMP, while the blue lines represent the settings without CodePMP.
+Figure 3: Best-of-N accuracy comparison: CodePMP-initialized models outperform baselines across various N values, showing superior ranking capabilities.
 4.2.2 BoN Accuracy Results
-We evaluate BoN accuracy across reasoning tasks with and without CodePMP initialization.
-As shown in Figure 8(a) and Table 8(b), RM finetuned with CodePMP initialization consistently achieves higher BoN accuracy across both mathematical and logical reasoning tasks for 1.5B and 7B models.
-This highlights CodePMP’s effectiveness in improving RM’s group-wise ranking performance.
-Across different values of N, RM models initialized with CodePMP maintain their lead, showing robust improvement even as N increases to 256.
-In contrast, RM without CodePMP shows a sharp decline in accuracy as N increases, underscoring the stability CodePMP provides, likely due to the diverse code-preference pairs used during training.
-For logical reasoning, the performance gap between CodePMP and non-PMP models is smaller, as $N$ is limited to 4, while in mathematical reasoning, $N$ reaches 256.
-This suggests that increasing $N$ further in logical reasoning could amplify the advantages of CodePMP in future evaluations.
-4.2.3 RM Sample Efficiency Comparison
-One key advantage of CodePMP is its ability to improve the sample efficiency of RM finetuning.
-To assess this, we conduct experiments with progressively larger sample sizes for RM finetuning.
-As indicated by (Kaplan et al., 2020), optimal results are achieved when the learning rate scheduler completes its decay at the end of training.
-Therefore, rather than evaluating intermediate checkpoints, we retrain models with varying sample sizes for optimal results.
-Figure 4 and  10 (Appendix E) show that as the sample size increases, RMs with CodePMP initialization consistently outperforms others in both BoN and RM accuracy.
-Notably, RMs finetuned with CodePMP initialization using just 0.5k samples surpasses RMs finetuned without CodePMP initialization using 40k samples on mathematical tasks, demonstrating CodePMP’s significant advantage in sample efficiency.
-However, as sample size increases, this advantage diminishes slightly, suggesting that with much larger datasets, CodePMP’s benefit may become less pronounced, but the cost of manual labeling remains a key consideration.
-4.2.4 The Importance of Scalable PMP
+Evaluations across reasoning tasks demonstrate that CodePMP-initialized RMs consistently achieve higher BoN accuracy on both mathematical and logical reasoning tasks for all model sizes (Figure 3).
+CodePMP models maintain performance advantages even as N increases to 256, while non-CodePMP models exhibit significant accuracy degradation at higher N values, highlighting CodePMP’s stability.
+This aligns with research on BoN sampling (Chow et al., 2024) that identifies an inflection point where performance typically deteriorates beyond certain N thresholds due to increased base policy stochasticity and verifier misalignment.
+CodePMP-initialized models demonstrate greater stability at higher N values, suggesting improved alignment with true reward signals and enhanced robustness to noise amplification inherent in large-N sampling.
+For logical reasoning, the performance gap appears smaller as testing was limited to N=4, while mathematical reasoning extended to N=256, suggesting potential for amplified advantages in logical reasoning with increased N values.
+4.2.3 Sample Efficiency Analysis
+To assess CodePMP’s impact on sample efficiency, we evaluated models with varying finetuning dataset sizes following best practices (Kaplan et al., 2020).
+Figure 4 shows that CodePMP-initialized models consistently outperform baselines across all dataset sizes, with CodePMP achieving with just 0.5k samples what baseline models require 40k samples to match—an 80× efficiency improvement.
+This advantage, while diminishing with larger datasets, significantly reduces annotation costs for developing effective reward models.
+4.2.4 Scalability Analysis
 A key benefit of using code data for PMP is the vast availability of publicly accessible, high-quality code-preference pairs, ensuring diversity.
 To validate scalability, we vary the number of training pairs for CodePMP and retrain models with different amounts of data.
-As shown in Figure 5, overall, increasing the number of code-preference pairs consistently improves BoN accuracy in both mathematical and logical reasoning tasks across model sizes, with no sign of diminishing returns.
+As shown in Figure 5, increasing the number of code-preference pairs consistently improves BoN accuracy in both mathematical and logical reasoning tasks across model sizes, with no sign of diminishing returns.
 This indicates that further scaling the code-preference data would likely yield additional performance gains, underscoring the importance of building a scalable PMP pipeline.
-(a) GSM8K / 1.5B
-(b) MATH / 1.5B
-(c) Reclor / 1.5B
-(d) LogiQA2.0 / 1.5B
-(e) GSM8K / 7B
-(f) MATH / 7B
-(g) Reclor / 7B
-(h) LogiQA2.0 / 7B
-Figure 5: Increasing the number of code-preference pairs consistently improves Best-of-N accuracy in both mathematical and logical reasoning tasks across model sizes, with no evident signs of diminishing returns.
-Note that the horizontal axis is scaled by $\sqrt{2}$, and the gray dashed line represents the results without CodePMP.
+(a) GSM8K / 7B
+(b) MATH / 7B
+(c) Reclor / 7B
+(d) LogiQA2.0 / 7B
+Figure 4: Sample efficiency comparison for 7B models: CodePMP-initialized reward models achieve higher Best-of-N accuracy with the equivalent sample sizes, showing better data efficiency.
+Horizontal axis scales by $\sqrt{2}$.
+Green: with CodePMP; Blue: without CodePMP.
+(a) GSM8K / 7B
+(b) MATH / 7B
+(c) Reclor / 7B
+(d) LogiQA2.0 / 7B
+Figure 5: Scaling analysis of CodePMP for 7B models: more code-preference pairs consistently improve Best-of-N accuracy across reasoning tasks without diminishing returns.
+Horizontal axis scales by $\sqrt{2}$; gray dashed lines show baseline performance without CodePMP.
 5 Ablation Studies
-In this section, we present a detailed analysis of CodePMP design.
-Unless otherwise stated, all experiments used the 1B model due to resource limitations and present the results of mathematical reasoning due to page limitation.
-Results of logical reasoning refers to Appendix E.2.
+This section presents a detailed analysis of CodePMP design.
+Unless otherwise stated, all experiments use the 1B model due to resource limitations and present the results of mathematical reasoning due to page limitation.
+More ablation studies refer to Appendix.
 5.1 Impact of Pair Construction
-Model-Generated Pairs Comparison
+GitHub-Sourced Pairs vs Web-Crawled
+We compare GitHub-sourced code with web-crawled data (Askell et al., 2021) from platforms such as StackExchange and Reddit.
+As shown in Figure 6(b), GitHub-sourced pairs (“Source Code”) consistently outperform those from web platforms (“Webpage”), particularly as the number of solutions (N) increases.
+Moreover, the performance improvement of GitHub-sourced pairs shows no sign of plateauing, highlighting the importance of diverse, high-quality source code in building a scalable PMP pipeline.
+Model Generated Data vs Human Data
 We compare various pair construction methods generated by different models.
@@
 Given that code execution can generate reliable outputs, future work will explore incorporating execution feedback to create more accurate preference pairs.
-Web-Crawled vs GitHub-Sourced Pairs
-We also compare GitHub-sourced code with web-crawled data( Askell et al.
-(2021)) from platforms such as StackExchange and Reddit.
-As shown in Figure 7(a), GitHub-sourced pairs (“Source Code”) consistently outperform those from web platforms (“Webpage”), particularly as the number of solutions (N) increases.
-Moreover, the performance improvement of GitHub-sourced pairs shows no sign of plateauing, highlighting the importance of diverse, high-quality source code in building a scalable PMP pipeline.
 (a) Different construction methods.
 (b) Different pair sources.
-Figure 6: Comparisons of BoN accuracy across different construction methods and pair sources.
-(a) With or without EOC token.
-(b) Using WSD or WCD learning rate schedulers.
-Figure 7: Comparisons of BoN accuracy for different settings: EOC token and lr schedulers.
-5.2 Impact of EOC Token
-Experiments by( Askell et al.
-(2021)) show that adding an end-of-context (EOC) token to each sequence significantly improves overall performance.
-To explore its impact in the context of CodePMP, we compared performance with and without the EOC token.
-As shown in Figure 7(a), the EOC setting (“w/ EOC”) consistently underperform the setting without EOC tokens (“w/o EOC”) across different test tasks, which is opposed to ( Askell et al.
-(2021)) We attribute this discrepancy to the different model, data and evaluation settings.
-5.3 Impact of Learning Rate Schedulers
-In the CodePMP experiments, we use the warmup-stable-decay (WSD) learning rate scheduler( Hu et al.
-(2024)), which can effectively reduce the time required for scaling related experiments.
-Previous studies mainly employ a learning rate schedule with linear warmup followed by cosine decay, known as warmup-cosine decay (WCD).
-We compare the performance of WSD and WCD, as shown in Table 7(b), both schedulers yield similar results.
-Thus, to improve computational efficiency, we adopt the WSD scheduler for all experiments.
+Figure 6: Comparison of BoN accuracy across construction methods and data sources, demonstrating benefits of model-based construction and GitHub code.
 (a) BoN accuracies on mathematical reasoning.
 (b) BoN (N=4) accuracies on logical reasoning.
-Figure 8: CodePMP improves sample efficiency and Best-of-N performance of the Gemma-2B reward models on reasoning tasks, highlighting its broad applicability across diverse LLM architectures.
-Table 2: Performance gains on coding RM and general RM (RMBench) evaluations show that CodePMP not only improves reasoning tasks but also generalizes well across various tasks.
-RMBench
-Model PMP Coding Summary Chat Chat Hard Safety Reasoning
-1.5B 0.6841 0.4154 0.4804 0.5351 0.3665 0.2751
-0.758 0.6126 0.9050 0.4364 0.3698 0.6041
-7B 0.6912 0.5839 0.4972 0.5022 0.5240 0.6804
-0.7619 0.7668 0.9413 0.5373 0.4906 0.9116
-5.4 Validating CodePMP on Other LLMs
-To further evaluate the generalizability of CodePMP, we validate its performance on the widely adopted Gemma-2B model (Team et al., 2024).
-As illustrated in Figure 8, the application of CodePMP results in significant performance gains in both mathemantical reasoning and logical reasoning evaluations.
-This not only underscores the robustness of CodePMP but also demonstrates its broad applicability in improving sample efficiency and overall performance across diverse LLM architectures.
-5.5 Performance on Coding and General RM Benchmarks
-We evaluate CodePMP on both code-specific and general reward modeling benchmarks.
-The CodeUltraFeedback_binarized test set serves as an in-domain evaluation, while RMBench provides an out-of-domain assessment.
-As shown in Table 2, models finetuned with CodePMP initialization consistently outperform those without CodePMP across various model sizes.
-These results demonstrate that CodePMP not only enhances performance in reasoning tasks but also generalizes well across a range of RM benchmarks.
+Figure 7: Cross-architecture performance comparison: CodePMP enhances reasoning performance across different model families (Gemma2 and Llama3.2), showing broad applicability.
+5.2 Impact of Loss Function
+CodePMP integrates both Reward Modeling (RM) and Language Modeling (LM) loss components.
+To evaluate their contributions, we conducted experiments comparing three configurations: RM loss only, LM loss only, and the combined approach.
+As shown in Table 2, the combined loss function consistently outperforms single-loss variants across all Best-of-N evaluation settings, with particularly notable improvements on the challenging MATH dataset.
+This empirical evidence indicates a complementary relationship where RM loss enhances preference ranking while LM loss preserves general language capabilities, collectively yielding more robust reward model performance.
+BoN RM Loss LM Loss RM + LM Loss
+GSM8K
+N=32 0.834 0.8317 0.8393
+N=64 0.8362 0.8271 0.8453
+N=128 0.8332 0.8309 0.8362
+N=256 0.8271 0.8226 0.8484
+MATH
+N=32 0.344 0.376 0.418
+N=64 0.358 0.376 0.424
+N=128 0.366 0.354 0.434
+N=256 0.362 0.372 0.41
+Table 2: Loss function comparison.
+5.3 Cross-Architecture Generalization
+To assess CodePMP’s generalization capabilities beyond the Qwen architecture family, we evaluated its effectiveness with Gemma2 and Llama3.2 as PMP/RM backbones on GSM8K, MATH, Reclor, and LogiQA-v2 benchmarks.
+As shown in Figure 7, CodePMP: (1) Consistently enhances reasoning performance across all model families, and (2) Improves robustness at larger $N$ values, mitigating performance degradation observed in non-initialized models.These results demonstrate that CodePMP generalizes effectively across diverse model architectures, suggesting broad applicability of the approach.
+5.4 Performance on Larger Backbone Model
+To investigate CodePMP’s performance on larger model scales, we applied the technique to Qwen2-72B. Table 3 presents results across mathematical and logical reasoning tasks.
+BoN w/o PMP w/ PMP w/o PMP w/ PMP
+GSM8K MATH
+N=1 0.7718 0.7718 0.298 0.298
+N=4 0.8453 0.8453 0.424 0.424
+N=32 0.8529 0.8628 0.488 0.500
+N=256 0.8249 0.8400 0.506 0.514
+BoN Reclor LogiQA-v2
+N=4 0.894 0.918 0.7117 0.7927
+Table 3: Performance comparison on reasoning tasks for Qwen2-72B with and without CodePMP initialization. Note that only $N=4$ was tested for Reclor and LogiQA-v2.
+Results show consistent improvements with CodePMP initialization across all benchmarks.
+Notably, performance gains increase with larger $N$ values on challenging tasks like MATH, indicating that CodePMP’s benefits scale effectively to larger model architectures.
+The significant improvement on logical reasoning tasks further demonstrates CodePMP’s scalability and broad applicability.
+5.5 Performance on More Powerful Generator
+To determine whether CodePMP maintains its effectiveness with more sophisticated generators, we conducted experiments with two advanced models: Qwen2-Math-7B-Instruct (specialized for mathematical reasoning) and Qwen2.5-32B-Instruct (a substantially larger general-purpose model).
+BoN GSM8K MATH
+w/o PMP w/ PMP w/o PMP w/ PMP
+N=4 0.8544 0.8931 0.690 0.724
+N=32 0.8446 0.8795 0.643 0.698
+N=256 0.8256 0.8590 0.614 0.690
+Table 4: BoN accuracy with specialized mathematical generator (Qwen2-Math-7B-Instruct).
+BoN GSM8K MATH
+w/o PMP w/ PMP w/o PMP w/ PMP
+N=4 0.9604 0.9688 0.798 0.820
+N=32 0.9573 0.9581 0.768 0.792
+N=256 0.9566 0.9634 0.752 0.798
+Table 5: BoN accuracy with large-scale generator (Qwen2.5-32B-Instruct).
+Tables 4 and 5 demonstrate that CodePMP’s benefits persist across different generator architectures.
+With the specialized Qwen2-Math-7B-Instruct (Table 4), we observe substantial improvements on both GSM8K and MATH.
+These gains remain consistent with the much larger Qwen2.5-32B-Instruct model (Table 5), despite it being significantly larger than both the preference pair generation models (less than 7B parameters) and the reward model itself (Qwen2-7B).
+These findings demonstrate that reward models trained on synthetic preference data from smaller models can effectively guide more powerful and specialized generators, confirming CodePMP’s robustness and cross-scale applicability.
+This is particularly significant as it suggests that relatively modest investments in reward model training can yield benefits even when deployed with state-of-the-art generation systems.
+5.6 Performance on General RM Benchmarks
+We further evaluate CodePMP on general reward modeling benchmarks (RMBench) to assess its applicability beyond reasoning tasks.
+RMBench provides an out-of-domain assessment covering various tasks including summarization, chat quality, and safety.
+As shown in Table 6, models fine-tuned with PMP consistently outperform those without PMP across various model sizes and tasks.
+Model PMP RMBench
+Summary Chat Chat Hard Safety Reasoning
+1.5B 0.4154 0.4804 0.5351 0.3665 0.2751
+0.6126 0.9050 0.4364 0.3698 0.6041
+7B 0.5839 0.4972 0.5022 0.5240 0.6804
+0.7668 0.9413 0.5373 0.4906 0.9116
+Table 6: Performance on RMBench shows that CodePMP generalizes well across various general LLM tasks.
+These results demonstrate that CodePMP enhances performance not only in reasoning and coding tasks but also improves generalization across a broad range of RM benchmarks.
+These findings provide compelling evidence for CodePMP’s broad applicability across multiple domains beyond the reasoning tasks that were our primary focus.
 6 Related Works
 Reward Modeling
-In the context of RLHF, reward models (RMs) have traditionally employed ranking models like Bradley-Terry and Plackett-Luce to represent human preferences (Bradley & Terry, 1952; Plackett, 1975; Cobbe et al., 2021b; Saunders et al., 2022; Lightman et al., 2023b; Wang et al., 2023b; Uesato et al., 2022; Luo et al., 2024; Yu et al., 2024; Stiennon et al., 2020; Nakano et al., 2021).
-More recently, probability-based approaches (Zhao et al., 2023; Jiang et al., 2023) have emerged, offering more precise predictions.
+In the context of RLHF, reward models (RMs) have traditionally employed ranking models like Bradley-Terry and Plackett-Luce to represent human preferences (Bradley and Terry, 1952; Plackett, 1975; Cobbe et al., 2021; Saunders et al., 2022; Lightman et al., 2023; Wang et al., 2023b; Uesato et al., 2022; Luo et al., 2024; Yu et al., 2024; Stiennon et al., 2020; Nakano et al., 2021).
+More recently, probability-based approaches have emerged, offering more precise predictions.
 Additionally, models such as Critique-out-Loud (Ankner et al., 2024) enhance RMs by integrating natural language feedback.
-Generative reward models (GRMs) (Yang et al., 2024b) further boost sample efficiency.
+Generative reward models (GRMs) further boost sample efficiency.
 Preference Modeling Pretraining (PMP) (Askell et al., 2021) introduces a novel pretraining phase, utilizing large-scale pairwise ranking data to enhance RM performance.
@@
 Code Training
-The inclusion of code in LLM pretraining has led to marked improvements in tasks such as commonsense reasoning (Madaan et al., 2022) and mathematical problem-solving (Liang et al., 2022; Shao et al., 2024; Yang et al., 2024a).
-Furthermore, code enhances general reasoning capabilities (Muennighoff et al., 2023; Fu & Khot, 2022; Ma et al., 2023).
+The inclusion of code in LLM pretraining has led to marked improvements in tasks such as commonsense reasoning (Madaan et al., 2022) and mathematical problem-solving (Liang et al., 2022; Shao et al., 2024; Yang et al., 2024).
+Furthermore, code enhances general reasoning capabilities (Muennighoff et al., 2023; Fu et al., 2022; Ma et al., 2023).
 Recent studies (Dong et al., 2023; Ma et al., 2023) indicate that incorporating code during supervised finetuning strengthens LLMs, particularly in complex decision-making tasks.
@@
 Improving reasoning capabilities in LLMs remains a significant challenge, with various advanced methods being proposed.
-Chain of Thought (CoT) prompting (Wei et al., 2022; Fu et al., 2023) improves reasoning by generating intermediate steps, while CoT combined with supervised finetuning (SFT) further enhances performance (Cobbe et al., 2021a; Liu et al., 2024; Yu et al., 2023).
+Chain of Thought (CoT) prompting (Wei et al., 2022; Fu et al., 2023) improves reasoning by generating intermediate steps, while CoT combined with supervised finetuning (SFT) further enhances performance (Cobbe et al., 2021; Liu et al., 2024; Yu et al., 2023).
 Other approaches focus on expanding inference time computation, such as problem decomposition (Zhou et al., 2022), search-based methods like MCTS (Xu, 2023), and using LLMs as verifiers (Huang et al., 2022; Luo et al., 2023).
-Reward models, including outcome-based (ORM) and process-based (PRM), have also shown success, with PRM delivering superior results (Lightman et al., 2023a; Wang et al., 2023b).
-Distinct from these approaches, CodePMP introduces a scalable preference model pretraining phase that can integrate seamlessly with all the aforementioned techniques.
-7 Conclusion & Future Works
-This paper introduces CodePMP, a scalable pretraining approach that leverages code-preference pairs to improve reasoning capabilities in large language models.
-Experimental results validate that CodePMP significantly improves sample efficiency and boosts performance on reasoning tasks.
-For future work, we aim to extend CodePMP in two key directions.
-CodePrMP will focus on utilizing compiler and interpreter verifiability to provide low-cost process supervision signals.
-GenPMP will explore how to improve sample efficiency and the performance of generative reward models by integrating code data.
+Reward models, including outcome-based (ORM) and process-based (PRM), have also shown success, with PRM delivering superior results (Lightman et al., 2023; Wang et al., 2023b).
+Encouragingly, CodePMP introduces a scalable preference model pretraining phase that can integrate seamlessly with all the aforementioned techniques.
+7 Conclusion and Future Works
+We propose CodePMP, a scalable preference model pretraining method that leverages synthetic code-preference pairs to boost reasoning in large language models.
+Experiments demonstrate that CodePMP markedly enhances both sample efficiency and performance across diverse reasoning tasks, validating the effectiveness of code-based preference pretraining.
+Future directions include CodePrMP, which will utilize compiler/interpreter verifiability for low-cost process supervision, and GenPMP, aimed at improving generative reward models through code-based pretraining.
+Limitations
+Our current implementation has several limitations.
+First, the synthetic preference pairs rely on models with predetermined parameter sizes, potentially missing nuanced preference signals that more sophisticated approaches might capture.
+While we demonstrate broad applicability across model families, architectural differences may affect performance in ways not fully explored in this work.
+Our reliance on GitHub data introduces potential biases stemming from the composition of public repositories.
+Additionally, our evaluation focuses primarily on mathematical and logical reasoning, leaving the method’s effectiveness for other reasoning modalities (e.g., commonsense or causal reasoning) less thoroughly examined.
+Future work should address these limitations to further enhance the generalizability and robustness of the approach.
+Ethics Statement
+CodePMP introduces several important ethical considerations.
+By enhancing LLMs’ reasoning capabilities, it could significantly impact decision-making systems that affect human lives, necessitating careful deployment and monitoring.
+While we utilize publicly available code, we recognize the importance of intellectual property rights and have focused on data with permissive licenses.
+Our approach reduces reliance on human annotation, potentially mitigating certain biases while possibly introducing others derived from the training data or model preferences.
+These trade-offs require ongoing evaluation and refinement to ensure fair and beneficial applications.
+As with any technology that enhances AI capabilities, responsible deployment with appropriate safeguards is essential.
 Appendix A Hyperparameters
-We outline key hyperparameters in this Table 3.
-In the tables, WSD refers to the warmup-stable-decay learning rate scheduler( (Hu et al., 2024)), which has the benefit of reducing the time required for scaling law experiments.
-Table 3: Hyperparameters for CodePMP training, mathematical reasoning RM finetuning, and logical reasoning RM finetuning.
-CodePMP Mathematical RM Logical RM
-HP 1.5B 7B 1.5B 7B 1.5B 7B
-epoch 1 1 1 1 1 1
-bs 1024 1024 64 64 64 64
-lr 3e-6 1e-6 1e-6 3e-7 1e-5 1e-5
-lr scheduler WSD WSD WCD WCD WCD WCD
-warmup ratio 0.03 0.03 0.03 0.03 0.25 0.25
-decay ratio 0.1 0.1 - - - -
-weight decay 0.1 0.1 0 0 0 0
-max length 1024 1024 1024 1024 1024 1024
-Appendix B Training Pipeline
-Figure B presents an overview of the complete training pipeline.
-The process begins with base language model (LM) pretraining on trillions of tokens from general text, followed by a preference model pretraining (PMP) phase using billions of tokens from code preference pairs.
-Finally, the model is finetuned on a smaller, more specialized dataset relevant to reasoning tasks, typically consisting of millions of tokens.
-Figure 9: An overview of the complete training pipeline.
-Appendix C CodePMP Dataset
-As shown in Table 4, the constructed CodePMP dataset consists of a total of 28 million files, accounting for 19 billion tokens across various languages.
-The dataset is primarily composed of Python files, with 20 million files and 13.1 billion tokens, followed by Notebook files, contributing 3 million files and 2.1 billion tokens, and other programming languages with 5 million files and 3.8 billion tokens.
-This diverse dataset supports the pretraining phase, aiding the model in generalizing across multiple reasoning tasks.
-The average length of chosen responses varies slightly between different languages.
-Python files exhibit an average response length of 170 tokens for chosen responses and 167 tokens for rejected responses, while Notebook files have slightly shorter average lengths of 158 and 155.5 tokens, respectively.
-Other languages show the highest average response lengths, with 213.2 tokens for chosen and 210 tokens for rejected responses.
-Overall, the average length of chosen responses across the dataset is 194.5 tokens, while the average length of rejected responses is 189.9 tokens, indicating minimal bias in the dataset based on response length.
-Table 4: Amount and average length statistics of CodePMP dataset.
-Amount Statistics Average Length
-Language Files (M) Tokens (B) Chosen Rejected
-Python 20 13.1 170.0 167.0
-Notebook 3 2.1 158.0 155.5
-Other Languages 5 3.8 213.2 210.0
-Total 28 19.0 194.5 189.9
-Appendix D RM Finetuning Dataset
-D.1 Mathematical Reasoning
-The RM finetuning for mathematical reasoning uses the MathShepherd dataset( (Wang et al., 2023b)), which contains 444k query-response samples, with some queries having multiple distinct responses.
-We divide the dataset into a 400k training set and a 44k test set.
-For RM finetuning, we construct preference pairs by selecting both correct and incorrect responses for the same query.
-To form the 4.3k test set, we combine one positive and negative sample for each query from the original test set.
-We also create two training sets of different sizes: MathShepherd-preference-800k and MathShepherd-preference-40k.
-The 800k training set is built by combining multiple positive and negative samples for each query in the original training set, resulting in 800k samples.
-In contrast, the 40k training set randomly selects one positive-negative pair for each query, totaling 40k samples.
-D.2 Logical Reasoning
-D.2.1 Reclor
-Reclor is a human-annotated reading comprehension reasoning dataset, where each sample consists of a passage, a question, and multiple options.
-To create preference pairs, we combine the correct and incorrect options for the same question.
-This process results in a total of 14.5k preference pairs, with 14k pairs used for training and 1.5k for testing, forming the Reclor-preference dataset.
-D.2.2 LogiQA2.0
-For logical reasoning finetuning scaling analysis, we synthesize a preference dataset based on the logical reasoning dataset LogiQA2.0.
-LogiQA2.0 is a reading comprehension benchmark requiring discrete reasoning over passages, with 96k crowdsourced, adversarially created questions.
-To answer correctly, models must resolve references (which may point to multiple locations in the input) and perform discrete operations like addition, counting, or sorting.
-We use four models (Qwen2-7B-Instruct [^1] https://huggingface.co/Qwen/Qwen2-7B-Instruct, Qwen2-72B-Instruct [^2] https://huggingface.co/Qwen/Qwen2-72B-Instruct, DeepSeek-V2-Chat [^3] https://huggingface.co/deepseek-ai/DeepSeek-V2-Chat, and DeepSeek-Coder-V2-Instruct [^4] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Instruct) to sample queries from the LogiQA2.0 dataset multiple times, with sampling $topp=1$, and $topk\in\{0,0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$.
-Answer correctness is annotated using DeepSeek-7B-Math-Compare-Answer [^5] https://huggingface.co/Tianqiao/DeepSeek-7B-Math-Compare-Answer.
-Correct and incorrect answers are combined to create preference pairs, resulting in 1,019k pairs, with 977k used for training and 42k for testing, forming the LogiQA2.0 preference dataset for RM scaling analysis.
-D.3 CodeUltraFeedback_binarized
-CodeUltraFeedback_binarized [^6] https://huggingface.co/datasets/coseal/CodeUltraFeedback_binarized is a preference dataset in the code domain, consisting of 9.5k preference pairs.
-We randomly split the dataset, using 90% of the samples for finetuning training and 10% for testing RM accuracy.
-Appendix E More Experimental Results
-Due to the length of the paper, we only report important experimental results in the experiment and ablation sections. In this section, we present complementary results in more detail.
-E.1 RM Sample Efficiency Comparison: RM accuracy results
-(a) MathShepherd-pair / 1.5B
-(b) Reclor-pair / 1.5B
-(c) LogiQA2.0-pair / 1.5B
-(d) MathShepherd-pair / 7B
-(e) Reclor-pair / 7B
-(f) LogiQA2.0-pair / 7B
-Figure 10: Comparison of sample efficiency of RM finetuning: Trends of RM accuracy with sample size increases.
-Note: The horizontal axis increases exponentially with $\sqrt{2}$.
-Therefore, the further the data intervals, the more RM data is effectively saved.
-We use different colors to highlight the results of different model sizes.
-We use RM accuracy evaluation to compare the sample efficiency of RM finetuning, as shown in Figure 10.
-Consistent with the conclusion from the main experiment, CodePMP improves sample efficiency.
-Under the same sample conditions, it consistently delivers stable improvements, making RM training more efficient.
-E.2 Ablation Results on logical Reasoning
-Table 5 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for different Model-generated pair construction methods for CodePMP. Overall, the $6.7B\&1.3B$ setting is the best among them.
-Table 6 shows the comparison of the BoN accuracy performance of ReClor and LogiQA2.0 using GitHub source code and web crawled code data to build CodePMP training pairs.
-The results verify that Github code is a better source for pair construction.
-Table 7 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for CodePMP models with and without EOC. In general, the setup without using EOC brings better results.
-Table 8 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for WCD and WSD lr scheduler settings. The results below show that WSD lr scheduler brings better results than WCD.
-Table 5: Comparison of ReClor and LogiQA2.0 BoN accuracies for different Model-generated pair construction methods for CodePMP. Overall, the $6.7B\&1.3B$ setting is the best among them.
-Test Set
-RM Finetune Constructions Method ReClor LogiQA2.0
-ReClor-pair Source Code & 1.3B 0.550 0.4464
-Source Code & 1.3B-Des-Clip 0.600 0.4522
-Source Code & 6.7B 0.578 0.4605
-1.3B & 1.3B-Des-Clip 0.572 0.4745
-6.7B & 1.3B-Des-Clip 0.564 0.4809
-6.7B & 1.3B 0.608 0.5032
-LogiQA2.0-pair Source Code & 1.3B 0.708 0.5217
-Source Code & 1.3B-Des-Clip 0.704 0.5268
-Source Code & 6.7B 0.714 0.5198
-1.3B & 1.3B-Des-Clip 0.738 0.5402
-6.7B & 1.3B-Des-Clip 0.742 0.5274
-6.7B & 1.3B 0.734 0.5415
-Table 6: Comparison of the BoN accuracy performance of ReClor and LogiQA2.0 using GitHub source code and web crawled code data to build CodePMP training pairs.
-The results below show that Github code is a better source for pair construction.
-Test Set
-RM Finetune Data Source ReClor LogiQA2.0
-ReClor-pair Webpage 0.574 0.4898
-Github 0.582 0.4981
-LogiQA2.0-pair Webpage 0.742 0.5293
-Github 0.752 0.5504
-Table 7: Comparison of ReClor and LogiQA2.0 BoN accuracies for CodePMP models with and without EOC. In general, the setup without using EOC brings better results.
-Test Set
-RM Finetune Model ReClor LogiQA2.0
-ReClor-pair w/o EOC 0.582 0.4981
-w/ EOC 0.596 0.4617
-LogiQA2.0-pair w/o EOC 0.752 0.5504
-w/ EOC 0.686 0.4809
-Table 8: Comparison of ReClor and LogiQA2.0 BoN accuracies for WCD and WSD lr scheduler settings. The results below show that WSD lr scheduler brings better results than WCD.
-Test Set
-RM Finetune Model ReClor LogiQA2.0
-ReClor-pair WCD 0.552 0.4828
-WSD 0.582 0.4981
-LogiQA2.0-pair WCD 0.748 0.5204
-WSD 0.752 0.5504
-Table 9: Cross-validation evaluation: ReClor and LogiQA2.0 BoN accuracies, black numbers are the results of cross-validation.
-In the cross-validation evaluation, CodePMP can still bring stable improvements on different test sets.
-Test Set
-RM Finetune Model PMP ReClor LogiQA2.0
-ReClor-pair 1.5B 0.534 0.4592
-0.582 0.4981
-7B 0.724 0.5453
-0.756 0.5835
-LogiQA2.0-pair 1.5B 0.710 0.5293
-0.752 0.5504
-7B 0.748 0.6371
-0.794 0.6779
-E.3 Cross-validations on logical Reasoning
-We further conduct cross-dataset validation on logical reasoning.
-Specifically, we evaluate the RM models finetuned on the ReClor-pairs dataset using the LogiQA2.0 test set.
-Similarly, we evaluate the RM models finetuned on the LogiQA2.0-pairs dataset using the ReClor test set.
-As shown in Table 9, CodePMP consistently improves RM evaluation performance, demonstrating that the enhancements CodePMP brings to RM training are robust and generalizable.
-Note that the BoN accuracies of the RM trained with LogiQA2.0-pair on the Reclor test set are higher than those of the RM trained directly on the Reclor-pair, because the LogiQA2.0-pair dataset is three times larger than the Reclor-pair dataset.
+We outline key hyperparameters in this section.
+In the tables, WSD refers to the warmup-stable-decay learning rate scheduler (Hu et al., 2024), which has the benefit of reducing the time required for scaling law experiments.
+Specifically, Table 7 lists the hyperparameters for CodePMP training, Table 8 details those for mathematical reasoning RM fine-tuning, Table 9 covers logical reasoning RM fine-tuning, and Table 10 presents the hyperparameters for BON generation.
+Hyperparameter Qwen2-1.5B Qwen2-7B
+epoch 1 1
+batch size 1024 1024
+learning rate 3e-6 1e-6
+lr scheduler WSD WSD
+warmup ratio 0.03 0.03
+decay ratio 0.1 0.1
+weight decay 0.1 0.1
+max length 1024 1024
+Table 7: CodePMP training hyperparameters.
+Hyperparameter Qwen2-1.5B Qwen2-7B
+epoch 1 1
+batch size 64 64
+learning rate 1e-6 3e-7
+lr scheduler WCD WCD
+warmup ratio 0.03 0.03
+weight decay 0 0
+max length 1024 1024
+Table 8: Mathematical reasoning RM finetuning hyperparameters.
+Hyperparameter Qwen2-1.5B Qwen2-7B
+epoch 1 1
+batch size 64 64
+learning rate 1e-5 1e-5
+lr scheduler WCD WCD
+warmup ratio 0.25 0.25
+weight decay 0 0
+max length 1024 1024
+Table 9: Logical reasoning RM finetuning hyperparameters.
+Hyperparameter MetaMath-Mistral-7B
+temperature 0.7
+top-p 1
+Table 10: Best-of-N generation hyperparameters.
+Appendix B CodePMP Dataset Statistics
+Table 11 presents the average token lengths of responses in the CodePMP dataset.
+The similar lengths between chosen and rejected responses (194.5 vs.
+189.9 tokens) ensure that response length does not introduce bias in the learning process.
+The dataset comprises 28 million files totaling 19 billion tokens, with Python (13.1B tokens), Jupyter Notebooks (2.1B tokens), and other languages (3.8B tokens) providing diverse coverage that facilitates model generalization.
+Language Chosen Rejected
+Python 170.0 167.0
+Notebook 158.0 155.5
+Other Languages 213.2 210.0
+Total 194.5 189.9
+Table 11: Average token lengths of responses in the CodePMP dataset by language category.
+(a) MATH-shepherd / 1.5B
+(b) MATH-shepherd / 7B
+(c) Reclor+Logiqa / 1.5B
+(d) Reclor+Logiqa / 7B
+Figure 8: Comparison of sample efficiency of RM fine-tuning: Trends of RM accuracy with sample size increases.
+(a) MATH-shepherd / 1.5B
+(b) MATH-shepherd / 7B
+(c) Reclor+Logiqa / 1.5B
+(d) Reclor+Logiqa / 7B
+Figure 9: Comparison of sample efficiency of RM fine-tuning: Trends of Multi-choice accuracy or Best-of-4 with sample size increases.
+Appendix C Further Comparisons and Cross-Domain Evaluations
+C.1 Comparison with Majority Voting
+We compare CodePMP with a majority-voting baseline under the same experimental setup on GSM8K and MATH.
+Table 12 shows that CodePMP outperforms majority voting, especially on more complex tasks like MATH.
+Method GSM8K MATH
+CodePMP 0.8484 0.41
+Majority Voting 0.8453 0.37
+Table 12: Comparison of CodePMP and majority voting on GSM8K and MATH.
+C.2 Sample Efficiency Improvements on Reclor and LogiQA
+We finetune the RM on preference pairs using only Reclor or LogiQA and then evaluate them on their respective test sets.
+As shown in Figures 8 and 9, PMP demonstrates a clear advantage in sample efficiency, reflected in both RM accuracy and Best-of-N evaluation.
+The results reveal that even with substantially fewer training samples, reward models initialized with CodePMP achieve comparable or better performance than models trained from scratch with many more samples, highlighting the significant sample efficiency benefits of our approach for logical reasoning tasks.
+C.3 Performance on Coding Tasks
+We evaluate CodePMP’s effectiveness on actual code generation tasks by conducting two types of evaluations: reward model accuracy assessment and code generation evaluation.
+First, we assess the reward model’s accuracy on the CodeUltraFeedback benchmark, which consists of preference pairs in the code domain.
+We fine-tuned Qwen2 models on the CodeUltraFeedback_binarized dataset (8.5k preference pairs), both with and without CodePMP initialization.
+Table 13 presents the accuracy results across model sizes.
+Model PMP CodeUltraFeedback
+Accuracy
+1.5B 0.6841
+0.758
+7B 0.6912
+0.7619
+Table 13: Performance on CodeUltraFeedback benchmark shows that CodePMP improves in-domain code reward modeling.
+As shown in Table 13, reward models initialized with CodePMP consistently outperform those without PMP initialization on the CodeUltraFeedback benchmark.
+For the 1.5B model, CodePMP initialization improves accuracy from 0.6841 to 0.758, while for the 7B model, accuracy increases from 0.6912 to 0.7619.
+These results demonstrate that CodePMP effectively enhances reward models’ ability to evaluate code quality.
+Beyond reward model evaluation, we also assess whether this improved evaluation capability translates to better code generation outcomes using the HumanEval benchmark.
+For this evaluation, we used deepseek-coder-6.7b-instruct as the generator and Qwen2-7B as the reward model (RM).
+We fine-tuned the RM on the same CodeUltraFeedback_binarized dataset, both with and without CodePMP initialization.
+Table 14 presents Pass@1 (0-shot) results under different $N$ values.
+BoN Qwen2-7B w/o PMP Qwen2-7B w/ PMP N=1 0.7134 0.7134 N=2 0.7317 0.7195 N=4 0.7073 0.7622 N=8 0.6890 0.7683 N=16 0.6951 0.7256 N=32 0.6585 0.7378 N=64 0.6829 0.7134 N=128 0.6707 0.7012 N=256 0.6707 0.7195
+Table 14: HumanEval results (Pass@1, 0-shot) for different numbers of sampled solutions $N$. The generator is deepseek-coder-6.7b-instruct.
+The results in Table 14 indicate that CodePMP initialization provides a generally more stable and higher-accuracy selection mechanism compared to direct training, especially as $N$ varies.
+For most values of $N$, the model with CodePMP initialization achieves better Pass@1 scores, with particularly notable improvements at $N=4$ (0.7622 vs.
+0.7073), $N=8$ (0.7683 vs.
+0.6890), and $N=32$ (0.7378 vs.
+0.6585).
+Without CodePMP, we observe performance degradation at higher $N$ values, while CodePMP-initialized models maintain more consistent performance.
+This finding is particularly significant since HumanEval evaluates actual code generation rather than just preference prediction, demonstrating that the benefits of CodePMP extend beyond improved preference modeling to better code generation outcomes.
+Summary
+These additional experiments demonstrate that CodePMP:
+  • Outperforms majority voting in both simpler (GSM8K) and more challenging (MATH) settings.
+  • Demonstrates significant sample efficiency improvements on logical reasoning tasks (Reclor and LogiQA), with models initialized with CodePMP achieving better performance with fewer training samples.
+  • Provides more stable and accurate code evaluation on HumanEval, showing benefits for practical code generation tasks.
+Thus, CodePMP provides a scalable and effective approach to improving large language models across different domains and tasks.
+Appendix D Comprehensive Data Diversity Analysis
+To validate the quality of our synthetic data, we conducted comprehensive diversity analyses using established methodologies from the research on synthetic text data generation (Zhu et al., 2024).
+These analyses aim to demonstrate that our synthetic data maintains sufficient diversity while effectively capturing the distributions present in human-generated data.
+D.1 N-gram Feature Distribution Analysis
+We mapped text n-gram features to fixed hash buckets (100 buckets) and analyzed their distribution patterns to measure lexical diversity.
+Figures 10 and 11 show the comparison between human-generated data and our synthetic data.
+Figure 10: Unigram distribution comparison (left: distribution for human data, right: distribution for synthetic data).
+Figure 11: Bigram distribution comparison (left: distribution for human data, right: distribution for synthetic data).
+Table 15 presents the density values for unigram and bigram distributions across different data sources.
+Data Source Unigram Bigram Human 134,538.40 133,538.41 Strong Model 97,653.69 96,653.70 Weak Model 93,691.69 92,691.70
+Table 15: N-gram density values for human and synthetic data.
+The distribution graphs show that, compared to human data, our synthetic data has more uniform n-gram distributions, without the concentration peaks common in synthetic data.
+The density values further quantify this advantage—synthetic text’s n-gram density values (Strong Model: 97,653.69, Weak Model: 93,691.69) are significantly lower than human text (134,538.40), demonstrating more balanced distribution across hash buckets.
+D.2 Embedding Space Visualization
+To further evaluate the semantic diversity of our synthetic data, we mapped semantic features of both human and synthetic data to a 2D space, as shown in Figure 12.
+Figure 12: Embedding space visualization (left: distribution for human data, right: distribution for synthetic data).
+Both synthetic and human data show wide and dispersed distributions in the embedding space with highly overlapping distribution ranges, indicating our synthetic data captures a similarly broad semantic space as human data.
+D.3 KL Divergence Analysis
+We quantified the distribution differences between synthetic and human data using KL divergence to evaluate how closely our synthetic data approximates natural distributions.
+Table 16 presents these results.
+N-gram Human Internal Strong Model Weak Model (Bootstrap) vs Human vs Human 1-gram 0.2502 0.4290 0.4631 2-gram 0.6904 1.3500 1.4281 3-gram 1.3012 2.5660 2.6693
+Table 16: KL divergence values comparing different data distributions.
+These results demonstrate that the distribution differences between our synthetic data and human data fall within acceptable ranges relative to internal human data variation.
+D.4 Comprehensive Validation of Synthetic Preference Data
+Our synthetic data generation approach relies on two key assumptions: (1) larger models from the same family produce higher-quality code than smaller ones, and (2) this quality difference creates consistent preference signals suitable for training.
+We conducted both theoretical and empirical validation to confirm these assumptions.
+D.4.1 Validation of Strong-Weak Model Ability Differences
+To validate our first assumption, we analyzed ability differences between strong and weak models across various code-related dimensions.
+(a) Same model family
+(b) Different model families
+Figure 13: Radar charts showing ability differences between strong and weak LLMs across various dimensions.
+Figure 13(a) provides strong evidence supporting our assumption: when using models from the same architectural family with different parameter counts, the stronger model consistently outperforms the weaker model across all ability dimensions.
+This uniform superiority ensures that synthetic preference pairs have clear and consistent quality differences, creating reliable signals for training preference models.
+For comparison, Figure 13(b) shows what happens when models from different families are paired.
+Here, we observe irregular and inconsistent differences, with some dimensions showing negligible gaps or even inversions.
+Such inconsistencies could potentially introduce noise into the preference signals, undermining training data quality.
+D.4.2 External Evaluation of Preference Consistency
+Having confirmed the underlying ability differences, we next validated whether these differences translate to consistent preference judgments.
+We conducted preference annotation experiments using GPT-4o to evaluate the consistency of preferences in our dataset.
+The results show that our synthetic CodePMP data achieved a preference consistency rate of 75.12%.
+This is notably higher than the more costly CodeUltraFeedback preference dataset (71.56%), demonstrating that the preference distinction in our synthetic data (based on our assumption that "larger models generate better code than smaller models") is sufficiently clear and consistent.
+This external validation reinforces our first finding - not only do larger models from the same family consistently outperform smaller ones across all dimensions, but this performance gap is readily detectable by strong evaluator models, resulting in consistent preference judgments.
+As CodePMP is fundamentally a pretraining process, we deliberately simplified our assumptions to enable scalable preference data creation with minimal additional validation.
+Our multi-faceted validation approach confirms that this simple yet effective methodology produces high-quality, consistent preference data suitable for large-scale pretraining.
+D.5 Source Data Quality and Experimental Validation
+Our method achieves excellent diversity due to the high quality of our source data:
+  • We collected over 130 million code snippets from GitHub, covering all common programming languages and task types on open-source platforms, ensuring breadth and depth in our source data.
+Furthermore, our experimental results validate the effectiveness of our synthetic data diversity:
+  • As shown in Figure 6(a), our synthesis strategy outperforms preference pairs constructed directly from source code, indirectly proving that our synthesis process enhances data diversity and quality.
+This comprehensive diversity analysis confirms that our synthetic data generation approach produces high-quality, diverse data that effectively captures the distribution characteristics of human-generated code.
+The balanced distribution patterns and semantic coverage demonstrate that our synthetic data is well-suited for training robust reward models.
+Appendix E Detailed Implementation of CodePMP
+In this section, we provide a detailed overview of the Code Preference Model Pretraining (CodePMP) implementation.
+The following description illustrates the systematic process of generating and utilizing code-preference pairs for pretraining preference models, which can then be fine-tuned for downstream reasoning tasks.
+The algorithm begins with a source code repository, a strong CodeLLM (in our implementation, deepseek-coder-6.7b-instruct), and a weaker CodeLLM (deepseek-coder-1.3b-instruct).
+First, descriptions are generated for each code snippet in the repository using the strong model.
+For each description, the strong model generates a high-quality chosen response, while the weaker model generates a less optimal rejected response.
+These pairs are used to calculate both language modeling loss (on the responses) and reward modeling loss (comparing chosen vs.
+rejected responses).
+The final training objective combines these two loss components.
+This scalable approach allows for creating millions of preference pairs without expensive human annotation, providing an effective initialization for reward models that will later be fine-tuned on specific reasoning tasks.
+Algorithm 1 Code Preference Model Pretraining
+0:  Source code repository $S$, Strong CodeLLM $M_{\text{strong}}$, Weak CodeLLM $M_{\text{weak}}$
+0:  Pretrained Model
+  Input: Source code $S$
+  Summarize description $D$ using $M_{\text{strong}}$ on $S$
+  for each $D_{i}\in D$ do
+     Generate Chosen Response using $M_{\text{strong}}$
+     Generate Rejected Response using $M_{\text{weak}}$
+  end for
+  Calculate LM Loss $\mathcal{L}_{\text{LM}}$ on Response
+  Calculate RM Loss $\mathcal{L}_{\text{RM}}$ using Chosen Response and Rejected Response
+  Train PMP Model using $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{RM}}+\mathcal{L}_{\text{LM}}$
 Appendix F Logical Reasoning Evaluation Examples
@@
 In principle, this method is equivalent to the Best-of-4 test.
-Table 10: Examples from the Reclor test set, which consists of multiple-choice questions based on a given passage.
-ID Text Options Answer
-12824 Mayor: Four years ago, when we reorganized the city police department in order to save money, critics claimed that the reorganization would make the police less responsive to citizens and would thus lead to more crime.
-The police have compiled theft statistics from the years following the reorganization that show that the critics were wrong.
-There was an overall decrease in reports of thefts of all kinds, including small thefts.
-Question: Which of the following, if true, most seriously challenges the mayor’s argument?
-1.
-In other cities where police departments have been similarly reorganized, the numbers of reported thefts have generally risen following reorganization.
-2.
-When city police are perceived as unresponsive, victims of theft are less likely to report thefts to the police.
-3.
-The mayor’s critics generally agree that police statistics concerning crime reports provide the most reliable available data on crime rates.
-4.
-The mayor’s reorganization of the police department failed to save as much money as it was intended to save.
-1
-218 Jupiter is a gas giant planet and the largest planet in the solar system.
-Its mass is 2.5 times the total mass of the other seven planets in the solar system.
-Observations have found that most of the more than 70 moons surrounding Jupiter are composed of water ice.
-Therefore, Jupiter’s atmosphere should contain a considerable amount of water.
-Question: Which of the followings, if true, can best support the above statement?
-1.
-After hundreds of millions of years, the satellite may slowly fall onto the planet.
-2.
-Many of the water in interstellar space exists in gaseous form.
-3.
-Uranus is also a gas giant planet, and it has been confirmed that it contains a lot of water ice.
-4.
-The satellite and the planets around it were formed from the same gas and dust at the same time.
+Table 17: Examples from the Reclor test set, illustrating multiple-choice questions format with passages and questions.
+ID Passage Question Ans.
+12824 Mayor: When we reorganized the police department, critics claimed it would make police less responsive and lead to more crime.
+Statistics show an overall decrease in thefts after reorganization.
+Which statement most challenges the mayor’s argument?
+(1) Similar reorganizations in other cities led to increased thefts.
+(2) Unresponsive police reduce theft reporting rates.
+(3) Critics agree police statistics are reliable.
+(4) The reorganization saved less money than planned.
+2
+218 Jupiter is the largest planet with mass 2.5 times that of all other planets combined.
+Most of Jupiter’s 70+ moons are water ice.
+What best supports that Jupiter’s atmosphere should contain water?
+(1) Satellites may eventually fall onto planets.
+(2) Interstellar water exists as gas.
+(3) Uranus, also a gas giant, contains water ice.
+(4) Satellites and planets form from the same materials.
 3
-10376 Lake Dali is a barrier lake on the plateau formed by volcanic eruptions.
-Like salmon living in the sea, Hua Zi fish-Leuciscus waleckii, which lives in a brackish lake, must migrate to the upper reaches of the Tanshui River to spawn and breed, although the four rivers currently flowing into Lake Dali are inland rivers, and none of them leads to the sea.
-Scientists are still convinced that the Huaziyu in Lake Dali first migrated from the ocean.
-Question: Which of the following options, if true, provides the best explanation for scientists’ beliefs?
-1.
-The Leuciscus waleckii that lives in the waters such as Heilongjiang is twice as big as the Leuciscus waleckii fish in Lake Dari.
-2.
-The caught Hua Zi fish can only survive for a day or two after being put into sea water or fresh water, and will decay quickly after death.
-3.
-Melting glaciers will form Lake Dali, and the overflowing lake was once connected to the Liao River, which flowed into the ocean.
-4.
-The researchers put the fry of Hua Zi fish in Dali Lake into Gainao thousands of miles away, and the culture was successful.
+10376 Lake Dali fish must migrate to river headwaters to breed, though no rivers connect to the sea.
+Scientists believe these fish originally came from the ocean.
+What best explains scientists’ belief?
+(1) Similar fish elsewhere are larger.
+(2) The fish quickly die in sea/fresh water.
+(3) Lake Dali was once connected to an ocean-bound river.
+(4) Fish from Lake Dali survived in far-away lakes.
 2
-13334 It is repeatedly claimed that the dumping of nuclear waste poses no threat to people living nearby.
-If this claim could be made with certainty, there would be no reason for not locating sites in areas of dense population.
-But the policy of dumping nuclear waste only in the more sparsely populated regions indicates, at the very least, some misgiving about safety on the part of those responsible for policy.
-Question: Which one of the following, if true, would most seriously weaken the argument?
-1.
-Until there is no shred of doubt that nuclear dumps are safe, it makes sense to situate them where they pose the least threat to the public.
-2.
-There are dangers associated with chemical waste, and it, too, is dumped away from areas of dense population.
-3.
-In the event of an accident, it is certain that fewer people would be harmed in a sparsely populated than in a densely populated area.
-4.
-Dumping of nuclear waste poses fewer economic and bureaucratic problems in sparsely populated than in densely populated areas.
+13334 If nuclear waste posed no threat, it could be placed in populated areas.
+But it is only dumped in sparsely populated regions, suggesting safety concerns.
+What would most weaken this argument?
+(1) Uncertain safety justifies minimal risk placement.
+(2) Chemical waste is also dumped away from population.
+(3) Accidents affect fewer people in sparsely populated areas.
+(4) Remote locations reduce bureaucratic complications.
 3